### v3

# OptiRefine: Densest subgraphs and maximum cuts with $k$ refinements

Abstract
Data-analysis tasks often involve an iterative process, which requires refining previous solutions. For instance, when analyzing social networks, we may obtain initial communities based on noisy metadata, and we want to improve them by adding influential nodes and removing non-important ones, without making too many changes. However, classic optimization algorithms, which typically find solutions from scratch, potentially return communities that are very dissimilar to the initial one. To mitigate these issues, we introduce the OptiRefine framework. The framework optimizes initial solutions by making a small number of refinements, thereby ensuring that the new solution remains close to the initial solution and simultaneously achieving a near-optimal solution for the optimization problem. We apply the OptiRefine framework to two classic graph-optimization problems: densest subgraph and maximum cut. For the densest-subgraph problem, we optimize a given subgraph’s density by adding or removing $k$ nodes. We show that this novel problem is a generalization of $k$-densest subgraph, and provide constant-factor approximation algorithms for $k=\Omega(n)$ refinements. We also study a version of maximum cut in which the goal is to improve a given cut. We provide connections to the maximum cut with cardinality constraints and provide an optimal approximation algorithm in most parameter regimes under the Unique Games Conjecture for $k=\Omega(n)$ refinements. We evaluate our theoretical methods and scalable heuristics on synthetic and real-world data and show that they are highly effective in practice.
1 Introduction
Graphs are commonly used to model entities and their relationships in various application domains. Numerous methods have been developed to address diverse graph mining tasks, including analyzing graph properties, finding patterns, discovering communities, and achieving other application-specific objectives. These graph mining tasks are generally formulated as optimization problems. Notable examples of such problems include graph-partitioning tasks, identifying densest subgraphs or determining the maximum cut of a graph.
In real-world applications, graph mining is often an iterative process that does not start from scratch. This process often begins with preliminary and suboptimal solutions, based on metadata or existing results. Subsequently the solution shall be refined and improved. Specifically, we seek to modify the initial solution $U$ into a new solution $U^{\prime}$ while ensuring that the solution $U^{\prime}$ must be close enough to $U$. More concretely, consider the following two examples:
  1. 1.
    Consider the task of detecting the community of users in an online social network who are interested in political commentary in a given country, say, Germany. Whether a user is related to political commentary in Germany can be obtained, for instance, based on metadata, such as the location data they are sharing and the interests they indicated when they joined the network. Let us call this set of users $U$. Some of those users, however, may not be active in the social network or changed their interests. Instead, there might be other users, unknown to us or from different countries, who are more active and also engage in political commentary regarding Germany. We could then seek to identify a set of users $U^{\prime}$, which is close enough to $U$ (so that the community is still on the same topic), and which has a high graph density (so that it consists of a set of influential users who are tightly connected with each other).
  2. 2.
    Next, suppose that a network host aims to maximize the diversity of the social network by altering a few users’ exposure to diverse news outlets (which can be realized, for instance, through recommendation). As formulated by recent work [Matakos et al., 2020], the exposure of each user can be modeled by a value in the set $\{-1,1\}$, e.g., representing two sides on a polarized topic. The network diversity can be measured by the size of the cut indicated by the partition $(U,\bar{U})$, where $U$ consists of the users whose exposure is 1-1 and $\bar{U}$ consists of the users whose exposure is $1$. Now the goal of increasing the network diversity corresponds to changing the exposure of a small set of users in order to maximize the cut. To ensure that not too many users are affected, the network host must ensure that the new assignment is close to the original one.
The above examples present a solution refinement process. Specifically, an initial solution is provided, such as an online community in the first example or the exposure of all network users to news outlets in the second example. The objective is to refine the solution so as to maximize an objective function, i.e., either finding a densest subgraph or a maximum cut, while maintaining proximity to the initial solution. The refinement process we consider in this paper appears in many graph-mining tasks. It naturally generalizes classic optimization algorithms, such as finding the densest subgraph or the maximum cut of a graph, by incorporating existing initial solutions.
From an algorithmic standpoint, standard algorithms designed for classical problems do not easily adapt to these new tasks that involve an initial solution and aim for the best refinement. To address this, we introduce the OptiRefine framework, which models this class of optimization problems. We apply our framework to graph-partitioning problems [Feige and Langberg, 2001], specifically focusing on the well-known densest subgraph and max-cut problems. We refer to our problems as graph-partition with $k$ refinements (GP-kkR), formally defined in Definition 7.
We apply our framework to two classic problems in data mining and combinatorial optimization: The densest-subgraph problem is commonly used to identify tightly connected groups of entities, with applications such as finding trending stories [Angel et al., 2012], detecting bots and fraudulent content in social media [Beutel et al., 2013], and identifying correlated genes in gene-expression data [Saha et al., 2010], among others. The max-cut problem has applications in graph-partitioning settings [Ding et al., 2001], in discovering conflicting communities in social networks [Bonchi et al., 2019], and in measuring the diversity of exposure to news outlets in online social networks [Matakos et al., 2020]. The methods we introduce in this paper are particularly useful when a community has been identified and there is a need to find another community that is close to the current one.
We identify a close relationship between (1) the problems in our OptiRefine framework and (2) the corresponding problems with cardinality constraints on the solution size. For both the densest subgraph and max-cut problems defined under the OptiRefine framework, we obtain approximation algorithms that nearly match the best approximation ratios of algorithms that solve the cardinality-constrained versions of these problems. We believe that our insights will be helpful in applying the OptiRefine framework to a broad range of problems and in developing practical algorithms with theoretical guarantees for a wide range of applications.
1.1 Our results
Refer to caption Refer to caption
(a) $k$-Densify (b) Max-Cut-kkR
Figure 1: The approximation ratios of the algorithms we present as a function of $k$. The approximation ratio for DS-kkR equals to the approximation ratio of $k$-Densify times $\frac{1-c}{1+c}$, where $c$ is such that $k\leq c\lvert U\rvert$. For $k$-Densify, we assume that the black-box solver is the algorithm for DkkS by Asahiro et al. [2000]. When $k=n/2$, the approximation ratio is at least $0.583$ with the SDP algorithm and at least $0.5$ with the black-box solver. For Max-Cut-kkR, we assume that the black-box solver is the algorithm for Max-Cut by Goemans and Williamson [1994] When $k=n/2$, the approximation ratio is at least $0.643$ with the SDP algorithm, at least $0.250$ with the black-box solver, and $0.858$ with the SoS algorithm.
First, we apply our OptiRefine framework to the classic densest subgraph and max-cut problems. For the densest-subgraph problem, we introduce the problem densest subgraph with $k$ refinements (DS-kkR) in which the input consists of a weighted graph $G=(V,E,w)$, an initial subset of vertices $U\subseteq V$ and an integer $k\in\mathbb{N}$. The goal is to add or remove $k$ vertices from $U$ to obtain a refined set of vertices $U^{\prime}$ that maximizes the density of its induced subgraph, denoted by $G[U^{\prime}]$ (see Problem 1 for the formal definition). For the max-cut problem, we introduce the problem of max-cut with $k$ refinements (Max-Cut-kkR) in which we are given a weighted graph $G=(V,E,w)$, an initial subset of vertices $U\subseteq V$ and an integer $k\in\mathbb{N}$. The goal is to find a set of vertices $U^{\prime}\subseteq V$ such that $U$ and $U^{\prime}$ differ by $k$ vertices and the cut $(U^{\prime},V\setminus U^{\prime})$ is maximized (see Problem 4 for the formal definition). For both DS-kkR and Max-Cut-kkR, we give reductions showing that they are closely connected to their corresponding classic problems, densest subgraph and maximum cut, respectively. In particular, we show that DS-kkR can be solved using solvers for the classic densest $k$-subgraph problem [Feige and Seltser, 1997] and the $k$-Densify problem [Matakos and Gionis, 2022]. Furthermore, Max-Cut-kkR can be solved using a solver for the classic max-cut problem [Goemans and Williamson, 1994] with only constant factor loss in approximation ratio. The approximation ratios using the black-box solvers are shown in Figure 1 with label Black-box. Moreover, we show that the reduction also works in the other way, and we present the hardness results for DS-kkR and Max-Cut-kkR.
Second, we apply our OptiRefine framework to the more general maximum graph-partitioning problem (MAX-GP[Feige and Langberg, 2001; Han et al., 2002; Halperin and Zwick, 2002] (see Problem 6 for formal definition). Through parameter settings [Han et al., 2002], MAX-GP encompasses various problems including densest-$k$ subgraph (DkkS), max-cut with cardinality constraints (Max-Cut-CC), max-uncut with cardinality constraints (Max-Uncut-CC), and vertex cover with cardinality constraints (VC-CC). MAX-GP can be solved approximately using semidefinite program-based (SDP-based) approaches.
We define the MAX-GP under the OptiRefine framework as graph partition with $k$ refinements (GP-kkR) (see Problem 7 for the formal definition). To solve GP-kkR, we adapt the SDP-based approaches for MAX-GP by incorporating the initial node statuses (inside or outside the initial subgraph) and designing appropriate constraints in the relaxed SDP program. We also propose a systematic yet simple method to determine the search space for suitable parameters, as introduced in Lemma C.4. Additionally, we introduce two pairs of parameters in the Max-Cut-kkR problem to handle the case distinctions arising in the rounding step. These schemes help us to obtain better approximation ratios.
For these problems, we establish constant-factor approximation ratios when $k=\Omega(n)$ refinements are allowed, where $n$ is the number of vertices in the graph. Our approximation ratios closely match the SDP-based approximation ratios for DkkS and Max-Cut-CC, which our problems generalize. Our approach extends the results of Matakos and Gionis [2022] for $k$-Densify(formally defined in Problem 2) by providing approximation algorithms for broader graph classes (see Section 2 for details). The approximation ratios achieved by our SDP-based approaches for DS-kkR and Max-Cut-kkR are shown in Figure 1 with label SDP.
Third, for Max-Cut-kkR, we provide approximation algorithms that almost match the best-known algorithms for max-cut with cardinality constraints. In particular, we use the sum-of-squares hierarchy (SoS) [Lasserre, 2002] to show that for $k=\Omega(n)$, Max-Cut-kkR admits the same approximation ratio as the state-of-the-art algorithm for cardinality-constrained max-cut [Raghavendra and Tan, 2012]. This approximation ratio is optimal for almost all parameter choices $k=\Omega(n)$, assuming the Unique Games Conjecture [Khot, 2002]. We present the approximation ratio for Max-Cut-kkR using SoS in Figure 1(b).
From a practical perspective, we implement our theoretical algorithms with provable guarantees, including SDP-based algorithms, black-box solver-based algorithms. For comparison, we also implement the greedy heuristics. We evaluate these algorithms and heuristics on both synthetic and real-world datasets. For DS-kkR, our algorithms with black-box solvers are scalable and increase the densities of the given subgraphs. For Max-Cut-kkR, our SDP-based approaches significantly outperform the methods by Matakos et al. [2020].
Structure of the paper. This paper is structured as follows. In Section 3, we define DS-kkR, and we illustrate its connections to $k$-Densify and DkkS. Moreover, we present a black-box solution for DS-kkR by applying a solver for DkkS. Section 4 presents Max-Cut-kkR, and we illustrate its connections to Max-Cut-CC. Moreover, we consider a black-box solution for Max-Cut-kkR by applying a solver for Max-Cut. In Section 5, we use a general problem, graph partitioning with $k$ refinements (GP-kkR), that captures both DS-kkR and Max-Cut-kkR. We then propose an SDP-based algorithm and obtain constant approximation results assuming $k\in\Omega(n)$. In Section 6, we describe a sum-of-squares algorithm to optimally solve Max-Cut-kkR under certain regimes of $k\in\Omega(n)$. Section 7 presents our experiments, in which we evaluate our algorithms for DS-kkR using multiple datasets. Moreover, in the appendix, we present more experimental evaluations for DS-kkR and Max-Cut-kkR, as well as all omitted proofs from the main text.
1.2 Notation
Throughout the paper, we let $G=(V,E,w)$ be an undirected graph with non-negative edge weights. We set $n=|V|$ and $m=|E|$. For a subset of vertices $U\subseteq V$, we write $E[U]$ to denote the set of edges with both endpoints in $U$, and $G[U]$ to denote the subgraph $(U,E[U])$. The density of a subgraph $G[U]$ is defined as $d(U)=\frac{{w(E[U])}}{|U|}$, where ${w(E[U])}=\sum_{(i,j)\in E[U]}w(i,j)$. We write $\overline{U}$ to denote the complement of $U$, i.e., $\overline{U}=V\setminus U$. Given a partition $(U,\overline{U})$ of $V$, we write $E[U,\overline{U}]$ to denote the set of edges that have one endpoint in $U$ and one in $\overline{U}$. We let $\mathrm{cut}\!\left(U\right)=\sum_{(i,j)\in E[U,\overline{U}]}w(i,j)$ denote the sum of edge weights over the cut $(U,\overline{U})$. For two sets $A$ and $B$ we denote their symmetric difference by $A\triangle B=(A\setminus B)\cup(B\setminus A)$. The operator $\log$ stands for $\log_{2}$.
2 Related work
The densest-subgraph problem has received considerable attention in the literature. Here, we discuss only the results that are most relevant to our work. We refer to the recent surveys by Lanciano et al. [2024] and Luo et al. [2023] for more discussion.
The problem of the densest subgraph with $k$ refinements (DS-kkR), as introduced in this paper, is a generalization of densest $k$-subgraph (DkkS[Kortsarz and Peleg, 1993]. Despite significant attention in the theory community, there is still a large gap between the best approximation algorithms and hardness results for DkkS [Asahiro et al., 2002; Barman, 2015; Bhaskara et al., 2010; Feige, 2002; Feige et al., 2001; Lee and Ghoshal, 2022; Manurangsi, 2018]. Asahiro et al. [2002] propose a greedy-peeling algorithm with an approximation ratio of $k/n$. Feige and Langberg [2001] introduce an approximation algorithm based on an SDP relaxation and hyperplane rounding, achieving a better approximation ratio than $\frac{k}{n}$. Han et al. [2002] build on this method by incorporating partition information into the hyperplane rounding, resulting in an improved approximation ratio. Notably, Feige and Seltser [1997] demonstrate that the integrality gap for standard SDP relaxation is $\mathcal{O}(n^{1/3}\log n)$ when $k=\Omega(n^{1/3})$, excluding the possibility of utilizing the standard SDP-based approach to achieve a better approximation ratio for all settings of $k$. The best current polynomial-time algorithm for DkkS provides an approximation guarantee of $\Omega(n^{-1/4-\varepsilon})$ for any constant $\varepsilon>0$ [Bhaskara et al., 2010], utilizing a more combinatorial approach compared to the SDP-based methods. Other strategies include those that leverage the structure of the adjacency matrix, as seen in [Papailiopoulos et al., 2014], and scalable methods for solving the relaxation of DkkS, as seen in [Lu et al., 2024].
The unconstrained version of the densest-subgraph problem has been used in practice to find communities in graphs. Goldberg [1984] has shown that this problem can be solved in polynomial time and Charikar [2000] provides a $2$-approximation algorithm. Besides the case of static graphs, efficient methods have been designed for dynamic [Bhattacharya et al., 2015; Sawlani and Wang, 2020] and streaming settings [McGregor et al., 2015].
Many variants of the densest-subgraph problem have been studied. For instance, Sozio and Gionis [2010] consider finding a densely connected subgraph that contains a set of query nodes. Dai et al. [2022]; Ye et al. [2024] consider finding a dense subgraph whose vertices are close to a set of reference nodes and contain a small set of anchored nodes. Tsourakakis et al. [2013] observe that, in practice, the densest subgraph is typically large and aim to find smaller, denser subgraphs by modifying the objective function. Another line of work aims at finding subgraphs that are dense for multiple graph snapshots [Charikar et al., 2018; Jethava and Beerenwinkel, 2015; Semertzidis et al., 2019].
Matakos and Gionis [2022] introduce the $k$-Densify problem, where the goal is to add $k$ vertices to a given subgraph to increase its density. They obtain $O(\sigma)$-approximation algorithms for graphs that have a $\sigma$-quasi-elimination ordering. This class of graphs includes, for instance, chordal graphs. However, general graphs may not have a small $\sigma$-quasi-elimination ordering. [^1] Their algorithm requires a predefined $\sigma$. In contrast, our algorithms from Section 3 provide constant-factor approximations for general graphs, but we need the assumption $k=\Omega(n)$. As our hardness results show, this assumption is necessary to obtain $O(1)$-approximation algorithms, assuming the Small Set Expansion Conjecture [Raghavendra and Steurer, 2010].
The problem of max-cut with $k$ refinements (Max-Cut-kkR) is a generalization of max-cut with cardinality constraints (Max-Cut-CC[Frieze and Jerrum, 1997; Papadimitriou and Yannakakis, 1991], in which the size of the cut is restricted to be equal to $k$, for a given $k\in[n]$. The most studied version of Max-Cut-CC is known as Max-Bisection, in which $k=n/2$ [Austrin et al., 2016; Feige and Langberg, 2001; Frieze and Jerrum, 1997; Manurangsi, 2019; Raghavendra and Tan, 2012]. The best polynomial-time approximation algorithm for Max-Cut-CC achieves an approximation ratio of approximately $0.858$ [Raghavendra and Tan, 2012]. This ratio is improved to approximately $0.8776$ for Max-Bisection [Austrin et al., 2016]. With respect to hardness of approximation for Max-Cut-CC, the best result is due to Austrin and Stanković [2019] who give hardness of approximation with $k=\tau n$ as a function of $\tau$, assuming the Unique Games Conjecture [Khot, 2002]. A matching approximation algorithm for $\tau\in(0,0.365)\cup(0.635,1)$ is given by Raghavendra and Tan [2012]. Our approximation algorithm for Max-Cut-kkR obtains the same approximation ratio. As we show that Max-Cut-kkR generalizes Max-Cut-CC, in the parameter setting above, our algorithms are also optimal under the Unique Games Conjecture [Khot, 2002].
The problem of graph partitioning with $k$ refinements (GP-kkR) is a generalization of maximum graph partitioning (MAX-GP[Feige and Langberg, 2001; Han et al., 2002; Halperin and Zwick, 2002], which generalizes several graph partitioning problems. Besides the two problems DkkS, Max-Cut-CC that we generalize in this paper, two other problems, namely, max-uncut with cardinality constraints (Max-Uncut-CC), and vertex cover with cardinality constraints (VC-CC) are also presented as applications of MAX-GP. All of the problems can be solved by SDP-based approaches, and constant approximation results are obtained assuming $k\in\Omega(n)$, where the concrete approximation ratios depend on the value of $k$.
As for picking the correct value of $k$ in practice, we are not aware of previous work covering this question for DkS or for Max-Cut-CC, and thus also not for our problems. We suggest the following two strategies for DSkR: (1) Our greedy algorithms can be adapted such that they keep on including/removing nodes as long as the density of the subgraph is increased. Once the density would decrease by a node addition/deletion, they stop. This makes the algorithm completely parameter-free, since $k$ is determined solely by the number of nodes whose addition/removal increases the subgraph’s density. Beyond our greedy algorithms, this approach can be mimicked by running our SDP-based or black-box methods for multiple values of k and picking the one with the highest density. (2) Another approach could be to only add vertices to the subgraph as long as at least a $\theta$-fraction of their degree is to vertices that were contained in the original subgraph, for some threshold $\theta$. This ensures that only vertices are added which are closely connected to vertices from the original subgraph. Similar approaches can also be used for Max-Cut-kR (by replacing the density objective with the cut objective).
Finally, we remark that independently and concurrently with our project, Fellows et al. [2023] and Grobler et al. [2023] propose reconfiguration frameworks that have similarities with our OptiRefine framework. In addition, Dalirrooyfard et al. [2024] study the $r$-move-$k$-partition problem, where the goal is to minimize the multiway cut among $k$ partitions by moving $r$ nodes’ positions. We note that this line of research focuses on developing fixed-parameter algorithms for different sets of problems. Hence, their results apply in the setting of small values of $k$, whereas we concentrate on large values of $k$. Therefore, the concrete results of these papers are incomparable with ours, but highlight the importance of studying refinement problems.
3 Densest subgraph with $k$ refinements
In this section, we formally define the problem of finding the densest subgraph with $k$ refinements (DS-kkR) and other relevant problems. We then study the relationship between these problems.
Problem 1 (Densest subgraph with $k$ refinements (DS-kkR)).
Given an undirected graph $G=(V,E,w)$, a subset $U\subseteq V$, and an integer $k\in\mathbb{N}$, DS-kkR seeks to find a subset of vertices $C\subseteq V$ with $|C|=k$ such that the density $d(U\triangle C)$ is maximized.
In the definition of Problem 1, the operation $\triangle$ denotes the symmetric difference of sets and $d(U\triangle C)$ is the density of the subgraph $G[U\triangle C]$, which is obtained by removing the vertices in $U\cap C$ and adding the vertices in $C\setminus U$. It is important to note that when $G[U]$ is already the densest subgraph, $G[U\triangle C]$ may have a lower density than $G[U]$ due to the constraint $|C|=k$.
We note that it may also be interesting to study the problem with the inequality constraint $\lvert C\rvert\leq k$. However, from a theoretical point of view, the results of Khuller and Saha [2009] and the proof of Lemma 3.1 imply that (up to constant factors) the inequality-constraint version of the problem is at least as hard to approximate as the equality-constraint version. Therefore, we study the equality-constraint version, because it simplifies the analysis. From a more practical point of view, we can solve the inequality-constraint version by running the equality-constraint version for multiple values of $k$.
Below, in Theorem 5.1 we show that we can obtain a constant factor approximation algorithm for DS-kkR. In this section, we focus on its computational complexity and its relationship with other classic problems. First, we introduce the $k$-Densify problem, which serves as an intermediate step in our analysis.
Problem 2 ($k$-Densify [Matakos and Gionis, 2022]).
Given an undirected graph $G=(V,E,w)$, a set $U\subseteq V$ and an integer $\in\mathbb{N}, kk$-Densify seeks to find a set of vertices $C\subseteq V$ with $\lvert C\rvert=k$ such that $w(E[U\triangle C])$} is maximized. 222 We note that in the definition given by Matakos and Gionis [2022], the objective is $w(E[U\cup C])$}. However, they implicitly require that $k\leq n-\lvert U\rvert$. Observe that in this case, the $k$ nodes are always selected from $V\setminus U$ in order to maximize the objective function. Hence, the maximum of $w(E[U\cup C])$} is equal to $w(E[U\triangle C])$}. In Problem 2, we adopt the objective function $w(E[U\triangle C])$} to make it consistent with DS-kkR.
Note that the $k$-Densify problem has the same cardinality constraint as DS-kkR, i.e., $|C|=k$. However, the objective function of $k$-Densify maximizes $w(E[U\triangle C])$}, i.e., it maximizes the weight of the edges in the subgraph. That is unlike DS-kkR which maximizes the density $d(U\triangle C)$. Thus, while in DS-kkR we can add and remove vertices from the given subgraph, in $k$-Densify we can only add vertices to the subgraph. This is because removing vertices from the subgraph always causes a decrease of the sum of the edge weights of the subgraph. However, in some cases removing the vertices from the subgraph increases the density of the subgraph. In Figure 2, we give a concrete example showing that DS-kkR may find subgraphs which are more community-like and have higher density.
Refer to caption Refer to caption Refer to caption
(a) The input subgraph $G[U]$. The outer nodes of $U$ are connected to exactly one node in $V\setminus U$.
(b) The optimal solution $U_{k\text{-}{{\small D}\textsc{ensify}}}$ for $k$-Densify. The density of the subgraph is $1.5$.
(c) The optimal solution $U_{\text{{\small DS}-}k{\small R}}$ for DS-kkR. The density of the subgraph is $2.5$.
Figure 2: An example illustrating the difference between the optimal solutions for $k$-Densify and DS-kkR. Here, we set $k=6$. We illustrate the subgraphs of the initial solution and of the optimal solutions using dotted circles. The nodes in red are those selected by the optimal solutions of the $k$-Densify and DS-kkR problems, respectively. Clearly, the solution found by DS-kkR is more community-like than the solution returned by $k$-Densify, and also yields a higher density.
Finally, we introduce the densest $k$-subgraph (DkkS).
Problem 3 (Densest $k$-subgraph (DkkS)).
Given an undirected graph $G=(V,E,w)$ and an integer $k\in\mathbb{N}$, DkkS seeks to find a set of vertices $U\subseteq V$ with $|U|=k$ such that the density $d(U)$ is maximized.
3.1 Relationships of the problems
We show the connection of the three problems above. In the subsequent analysis, the approximation ratios that we derive will typically depend on $k$. To emphasize the dependence on $k$, we will denote the approximation ratios using the notation $\alpha_{k}$.
Relationship of DS-kkR and DkkS. First, we show that DS-kkR generalizes the classic densest $k$-subgraph (DkkS) problem.
Lemma 3.1.
Let $k\in\mathbb{N}$. If there exists an $\alpha_{k}$-approximation algorithm for DS-kkR with running time $T(n,m,k)$, then there exists an $\alpha_{k}$-approximation algorithm for DkkS running in time $\mathcal{O}(T(n,m,k))$.
Lemma 3.1 allows us to obtain hardness-of-approximation results for DS-kkR based on the hardness of DkkS [Lee and Ghoshal, 2022].
Corollary 3.1 (Hardness of approximation for DS-kkR).
Any hardness of factor $\beta_{k}$ for approximating DkkS implies hardness of approximating DS-kkR within a factor $\beta_{k}$. In particular, if $k=\tau n$ for $\tau\in(0,1)$, then DS-kkR is hard to approximate within $C\tau\log(1/\tau)$, where $C\in\mathbb{R}_{+}$ is some fixed constant, assuming the Small Set Expansion Conjecture [Raghavendra and Steurer, 2010].
Corollary 3.1 implies that for $k=o(n)$ (and, therefore, $\tau=o(1)$), it is not possible to obtain constant-factor approximation algorithms for DS-kkR. Additionally, for $k=\Omega(n)$, the problem is hard to approximate within a constant factor.
Relationship of $k$-Densify and DS-kkR. We show that DS-kkR can be effectively solved by applying an algorithm for $k$-Densify on the same input instance, with little impact on the approximation ratio. Specifically, Lemma 3.2 shows that running an $\alpha_{k}$-approximation algorithm for $k$-Densify on an instance of DS-kkR only loses a factor of factor $\frac{1-c}{1+c}$ in approximation, where $c\in(0,1)$ is a constant such that $k\leq c\lvert U\rvert$. We note that in practice, it is realistic that $c$ is small since we only want to make a small fraction of changes to $U$. This implies that $\frac{1-c}{1+c}$ is large; for instance, $\frac{1-c}{1+c}\geq 0.9$ if $c=5\%$. Notice that the condition $k\leq n-\lvert U\rvert$ is an implicit requirement from the definition of the $k$-Densify problem.
Lemma 3.2.
Let $(G=(V,E,w),U,k)$ be an instance for DS-kkR. Let us assume that $k\leq c\lvert U\rvert$ for some $c\in(0,1)$, and $k\leq n-\lvert U\rvert$. Applying an $\alpha_{k}$-approximation algorithm for $k$-Densify on $G$ provides a $\left(\frac{1-c}{1+c}\,\alpha_{k}\right)$-approximate solution for DS-kkR.
Relationship of $k$-Densify and DkkS. We show that an algorithm for the Densest $(k+1)$-Subgraph problem (D(k+1)(k\!+\!1)S) can be effectively utilized to solve $k$-Densify, with only small losses in the approximation ratio.
Lemma 3.3.
Suppose there exists an $\alpha_{k+1}$-approximation algorithm for D(k+1)(k\!+\!1)S, then there exists a $\frac{k-1}{k+1}\alpha_{k+1}$-approximation algorithm for $k$-Densify.
Proof sketch.
On a high level, the reduction from $k$-Densify to D(k+1)(k\!+\!1)S works as follows (see Appendix A.5 for details). Given an instance $(G=(V,E,w),U,k)$ for $k$-Densify, we build a graph $G^{\prime}$ in which we contract all vertices in $U$ into a special vertex $i^{*}$. Then we solve D(k+1)(k\!+\!1)S on the graph $G^{\prime}$. We then consider two cases: either $i^{*}$ is included in the (approximate) solution $U^{\prime}$ of D(k+1)(k\!+\!1)S, or not. If $i^{*}$ is part of the solution, we have added exactly $k$ vertices to $U$. Otherwise, we add all vertices from $U^{\prime}$ to $U$, except the vertex from $U^{\prime}$ of lowest degree in $G[U\cup U^{\prime}]$; this implies that we also added $k$ vertices to $U$. Then we can show that the $k$ additional vertices chosen by the D(k+1)(k\!+\!1)S-algorithm form a solution for $k$-Densify subgraph with the desired approximation guarantee. ∎
3.2 Black-box reduction
By combining the results from the previous section, we can now show that any approximation algorithm for DkkS implies (under some assumptions on $|U|$ and $k$) an approximation algorithm for DS-kkR. This result can be viewed as the reverse direction of the reduction in Lemma 3.1, and is obtained by applying Lemmas 3.2 and 3.3. Our formal result is as follows.
Theorem 3.4.
Suppose there exists an $\alpha_{k}$-approximation algorithm for the DkkS problem with running time $T(m,n,k)$. If $k\leq c\lvert U\rvert$, for some $c\in(0,1)$, there exists a $\left(\frac{1-c}{(1+c)}\,\frac{k-1}{k+1}\,\alpha_{k+1}\right)$-approximation algorithm for the DS-kkR problem with running time $\mathcal{O}(T(m,n,k+1)+m)$.
By utilizing the approximation algorithm for DkkS by Feige and Langberg [2001] in Theorem 3.4, we obtain Corollary 3.2. We include the assumption $k=\Omega(n)$ in the corollary, as it is a requirement of Feige and Langberg [2001].
Corollary 3.2.
If $\lvert U\rvert=\Omega(n)$, $k=\Omega(n)$, and $k\leq c\lvert U\rvert$ for a constant $c\in(0,1)$, there exists an $\mathcal{O}(1)$-approximation algorithm for DS-kkR.
In our experiments, we will implement the algorithm from Corollary 3.2 and denote it as black-box algorithm, since the algorithm is using a black-box solver for DkkS.
4 Max-cut with $k$ refinements
In this section, we study max-cut with $k$ refinements (Max-Cut-kkR). We start by providing formal definitions for Max-Cut-kkR and for cardinality-constrained max-cut (Max-Cut-CC). We then present a hardness result for Max-Cut-kkR through Max-Cut-CC. In the reverse direction, we show that any approximation algorithm for unconstrained Max-Cut provides an approximation algorithm for Max-Cut-CC, in a black-box reduction. Later, in Sections 5 and 6, we will also provide constant-factor approximation algorithms for Max-Cut-kkR.
We start by formally defining Max-Cut-kkR and give an example in Figure 3.
Problem 4 (Max-cut with $k$ refinements (Max-Cut-kkR)).
Given an undirected graph $G=(V,E,w)$, a subset $U\subseteq V$, and an integer $k\in\mathbb{N}$, Max-Cut-kkR seeks to find a set of vertices $C\subseteq V$ such that $|C|=k$ and $\mathrm{cut}\!\left(U\triangle C\right)$ is maximized.
As with the DS-kkR problem, observe that if $\mathrm{cut}\!\left(U\right)$ is the maximum cut, $\mathrm{cut}\!\left(U\triangle C\right)$ may have a lower value than $\mathrm{cut}\!\left(U\right)$ due to the equality constraint $|C|=k$.
(a) The initial partition of the graph. The cut value is $3$.
(b) The optimal solution when $k=1$. The cut value is $5$.
(c) The optimal solution when $k=6$. The cut value is $5$.
Figure 3: An illustration of Max-Cut-kkR on a graph with 77 nodes. The dotted line indicates the partition of the graph. The nodes with red color are the nodes selected by the optimal solution of the Max-Cut-kkR problem. The red and blue edges indicate that the edge is moved into or out of the cut set, respectively.
We also introduce the problem of maximum cut with cardinality constraints (Max-Cut-CC), which we use for the analysis of the hardness of Max-Cut-kkR.
Problem 5 (Max-cut with cardinality constraints (Max-Cut-CC)).
Given an undirected graph $G=(V,E,w)$, and an integer $k\in\mathbb{N}$, Max-Cut-CC seeks to find a set of vertices $U\subseteq V$ of size $|U|=k$ such that $\mathrm{cut}\!\left(U\right)$ is maximized.
Relationship of Max-Cut-kkR and Max-Cut-CC. First, we show that Max-Cut-kkR is a generalization of Max-Cut-CC in Lemma 4.1.
Lemma 4.1.
Let $k\in\mathbb{N}$. If there exists an $\alpha_{k}$-approximation algorithm for Max-Cut-kkR with running time $T(n,m,k)$, then there exists an $\alpha_{k}$-approximation algorithm for Max-Cut-CC with running time $\mathcal{O}(T(n,m,k))$.
Combining Lemma 4.1 and the results of Austrin and Stanković [2019], we obtain the following hardness result for Max-Cut-kkR.
Corollary 4.1.
Any hardness of factor $\beta_{k}$ for approximating Max-Cut-CC implies hardness of approximating Max-Cut-kkR within factor $\beta_{k}$. In particular, suppose that $k=\tau n$ for $\tau\in(0,1)$. For $\tau\leq 0.365$ or $\tau\geq 0.635$, we have that $\beta_{k}\approx 0.8582$ and for $\tau\in(0.365,0.635)$ we have $\beta_{k}\in[0.8582,0.8786]$.
4.1 Black-box reduction
Finally, we show that algorithms for the (unconstrained) Max-Cut problem can be used to solve Max-Cut-kkR.
Theorem 4.2.
Consider an $\alpha$-approximation algorithm for Max-Cut with running time $T(m,n)$. There exists an algorithm for Max-Cut-kkR with running time $\mathcal{O}(T(m,n)+m)$. The approximation ratio when $k\leq n/2$ is $\alpha\left(1-\Theta\!\left(\frac{1}{n}\right)\right)\min\left\{4\frac{k^{2}}{n^{2}},\left(1-\frac{k}{n}\right)^{2}\right\}$. For $k\geq n/2$, the approximation ratio is $\alpha\left(1-\Theta\!\left(\frac{1}{n}\right)\right)\min\left\{4\frac{(n-k)^{2}}{n^{2}},\left(\frac{k}{n}\right)^{2}\right\}$.
In our reduction, we run an algorithm (e.g., the SDP relaxation by Goemans and Williamson [1994]) for Max-Cut without cardinality constraint, i.e., the cut may contain an arbitrary number of vertices on each side. Suppose this algorithm returns a cut $(U^{\prime},\overline{U^{\prime}})$. Then we make refinements to $U^{\prime}$ to bring it “closer” to $U$, i.e., we greedily move vertices until $U^{\prime}=U\triangle C$ for some set $C$ with $|C|=k$. Using the assumptions from the theorem, we obtain an $\mathcal{O}(1)$-approximation algorithm assuming that $k=\Omega(n)$. The pseudocode for this method is presented in the appendix.
5 Graph partitioning with $k$ refinements
The maximum graph-partitioning problem (MAX-GP) characterizes DkkS and Max-Cut-CC, as well as max-uncut with cardinality constraints (Max-Uncut-CC), and vertex cover with cardinality constraints (VC-CC) with specific parameter settings. Previous papers [Feige and Langberg, 2001; Han et al., 2002; Halperin and Zwick, 2002] provided SDP-based approximation algorithms for MAX-GP. Let us define MAX-GP the same way as [Han et al., 2002], by introducing parameters $c_{0},c_{1},c_{2},c_{3}$ into the MAX-GP objective function, specified in Table 1. Our analysis in the remainder of this section holds for these specific settings.
Table 1: Partition measures and corresponding values of $c_{0},c_{1},c_{2},c_{3}$ for the problems we study in this paper.
Problem GP-kkR Problem MAX-GP Measure $\mathcal{G}$ $c_{0}$ $c_{1}$ $c_{2}$ $c_{3}$
$k$-Densify DkkS $w(E[\cdot])$} $\frac{1}{4}$ $\frac{1}{4}$ $\frac{1}{4}$ $\frac{1}{4}$
Max-Cut-kkR Max-Cut-$C$ $\mathrm{cut}(\cdot)$ $\frac{1}{2}$ 0 0 $\frac{1}{2}$
Max-Uncut-kkR Max-Uncut-$C$ ${w(E[V])}-\mathrm{cut}(\cdot)$ $\frac{1}{2}$ 0 0 $\frac{1}{2}$
VC-kkR VC-$C$ ${w(E[\cdot])}+\mathrm{cut}(\cdot)$ $\frac{3}{4}$ $\frac{1}{4}$ $\frac{1}{4}$ $\frac{1}{4}$
Problem 6 (Maximum graph-partitioning (MAX-GP)).
Given an undirected graph $G=(V,E,w)$ where $\lvert V\rvert=n$ and an integer $k\in\mathbb{N}$, the predefined parameters $c_{0}$, $c_{1}$, $c_{2}$, $c_{3}$ and the partition measure $\mathcal{G}\!\left(\cdot\right)$, MAX-GP seeks to find a set of vertices $U\subseteq V$ where $|U|=k$, such that the partition measure $\mathcal{G}\!\left(U\right)$ is maximized.
In particular, let $\bm{x}\in\{-1,1\}^{n}$ be an indicator vector such that $x_{i}=1$ if and only if $i\in U$. MAX-GP is formulated as follows:
$\max_{\bm{x}}$ $\sum_{i<j}w(i,j)\,(c_{0}+c_{1}x_{i}+c_{2}x_{j}+c_{3}x_{i}x_{j}),$ (1)
such that $\sum_{i}x_{i}=-n+2k\text{ and }$
$\bm{x}\in\{-1,1\}^{n}.$
The predefined parameters and partition measures are presented in Table 1.
In this section, we generalize MAX-GP to take into account refinements of existing solutions. We call our generalized problem graph partitioning with $k$ refinements (GP-kkR) and define it formally below. The $k$-Densify and Max-Cut-kkR from above are instances of GP-kkR by specifying the partition measure $\mathcal{G}\!\left(\cdot\right)$.
Problem 7 (Graph partitioning with $k$ refinements (GP-kkR)).
Given an undirected graph $G=(V,E,w)$ where $\lvert V\rvert=n$, a subset $U\subseteq V$, and an integer $k\in\mathbb{N}$, the predefined parameters $c_{0}$, $c_{1}$, $c_{2}$, $c_{3}$ and the partition measure $\mathcal{G}\!\left(\cdot\right)$, GP-kkR seeks to find a subset of vertices $C\subseteq V$ with $\lvert C\rvert=k$, such that the partition measure $\mathcal{G}\!\left(U\triangle C\right)$ is maximized.
In particular, let $\bm{x}^{0}\in\{-1,1\}^{n}$ be an indicator vector such that $x^{0}_{i}=1$ if and only if $i\in U$ and let $\bm{x}\in\{-1,1\}^{n}$ be an indicator vector such that $x_{i}=1$ if and only if $i\in U\triangle C$. GP-kkR is formulated as follows:
$\max_{\bm{x}}$ $\sum_{i<j}w(i,j)\,(c_{0}+c_{1}x_{i}+c_{2}x_{j}+c_{3}x_{i}x_{j}),$ (2)
such that $\sum_{i}x^{0}_{i}x_{i}=n-2k\text{ and }$
$\bm{x}\in\{-1,1\}^{n}.$
The predefined parameters and partition measures are presented in Table 1.
Notice that when $U=\emptyset$ (and $\bm{x}^{0}=-\bm{1}$), the formulation of GP-kkR becomes the same as MAX-GP. Taking into consideration the initial solution ($U\neq\emptyset$) changes the formulation of the constraint part. Hence, we mainly illustrate the constraint. We use the indicator vector $\bm{x}^{0}\in\{-1,1\}^{V}$ to encode the set $U$, i.e., we set $x^{0}_{i}=1$ if $i\in U$ and $x^{0}_{i}=-1$ if $i\in V\setminus U$. We then set our solution $C$ to all $i$ such that $x_{i}\neq x^{0}_{i}$, i.e., $C=\{i\colon x_{i}\neq x^{0}_{i}\}$. Observe that the constraint of the integer program implies that $C$ contains exactly $k$ vertices: we have $x^{0}_{i}x_{i}=1$ if and only if $x^{0}_{i}=x_{i}$ and $x^{0}_{i}x_{i}=-1$ if and only if $x^{0}_{i}\neq x_{i}$. Hence, the sum of all $x^{0}_{i}x_{i}$ is equal to $n-2\ell$, where $\ell=|\{i\colon x_{i}\neq x^{0}_{i}\}|$ is the number of entries in which $\bm{x}$ and $\bm{x}^{0}$ differ. Since in the constraint we set $\sum_{i}x^{0}_{i}x_{i}=n-2k$, we get $|C|=k$.
The objective functions of GP-kkR are essentially the same as MAX-GP, as well as the parameter settings. Next, we illustrate the settings of $k$-Densify and Max-Cut-kkR. For Max-Uncut-kkR and VC-kkR, we refer readers to Han et al. [2002] for more details. For $k$-Densify, the term $(\frac{1}{4}+\frac{1}{4}x_{i}+\frac{1}{4}x_{j}+\frac{1}{4}x_{i}x_{j})$ equals $1$ if $x_{i}=x_{j}=1$ and equals 0 otherwise. Hence, the objective aims to maximize the number of edges in $G[U\triangle C]$, i.e., we maximize $w(E[U\triangle C])$}. For Max-Cut-kkR, the terms $(\frac{1}{2}-\frac{1}{2}x_{i}x_{j})=1$ if and only if $x_{i}\neq x_{j}$, i.e., when $i$ and $j$ are on different sides of the cut, otherwise $\frac{1}{2}-\frac{1}{2}x_{i}x_{j}=0$. Thus, the objective function sums over all edges with one endpoint in $U\triangle C$ and one endpoint in $\overline{U}\triangle C$, which is identical to $\mathrm{cut}\!\left(U\triangle C\right)$.
5.1 SDP-based algorithm
Next, we present our SDP-based algorithm. We will show that the SDP-based algorithm, as applied by Frieze and Jerrum [1997] to solve Max-Cut-CC and by Feige and Langberg [2001] to solve DkkS, can be adapted to solve our problem GP-kkR.
The semidefinite program. We state the SDP relaxation of the integer program (2) as follows, where $\bm{v}_{i}\cdot\bm{v}_{j}$ denotes the inner product of $\bm{v}_{i}$ and $\bm{v}_{j}$:
$\max_{\bm{v}_{0},\ldots,\bm{v}_{n}\in\mathbb{R}^{n}}$ $\sum_{i<j}w(i,j)\,(c_{0}+c_{1}\bm{v}_{0}\cdot\bm{v}_{i}+c_{2}\bm{v}_{0}\cdot\bm{v}_{j}+c_{3}\bm{v}_{i}\cdot\bm{v}_{j}),$ (3)
such that $\sum_{i}x_{i}^{0}\bm{v}_{i}\cdot\bm{v}_{0}=n-2k,$
$\sum_{i,j}x_{i}^{0}x_{j}^{0}\bm{v}_{i}\cdot\bm{v}_{j}=(2k-n)^{2},\text{ and }$
$\bm{v}_{i}\in\mathcal{S}_{n},\text{ for }i=0,\ldots,n.$
In the SDP, we replace each entry $x_{i}$ with the inner product $\bm{v}_{i}\cdot\bm{v}_{0}$ and each $x_{i}x_{j}$ with the inner product $\bm{v}_{i}\cdot\bm{v}_{j}$. All $\bm{v}_{i}$ are vectors from the unit sphere $\mathcal{S}_{n}$. For technical reasons, we add the constraint $\smash{\sum_{i,j}x_{i}^{0}x_{j}^{0}\bm{v}_{i}\cdot\bm{v}_{j}}=(2k-n)^{2}$, which is a relaxation of the constraint $\smash{(\sum_{i,j}x_{i}^{0}x_{i})^{2}}=(2k-n)^{2}$. This constraint will help us to bound the variance of the random-rounding procedure.
The algorithm. Our algorithm for GP-kkR consists of the following three steps:
  1. 1.
    We solve the SDP (3) to obtain a solution $\bm{v}_{0},\ldots,\bm{v}_{n}$.
  2. 2.
    We obtain an indicator vector $\overline{\bm{x}}\in\{-1,1\}^{n}$ (which we will later use to obtain $C$) as follows. We sample a unit vector $\bm{r}$ from $\mathcal{S}_{n}$ uniformly at random and obtain the indicator vector $\overline{\bm{x}}\in\{-1,1\}^{n}$ by setting $\overline{x}_{i}=1$ if $(\bm{r}\cdot\bm{v}_{0})(\bm{r}\cdot\bm{v}_{i})\geq 0$ and $\overline{x}_{i}=-1$ otherwise. We call this step hyperplane rounding.
  3. 3.
    As before, we set $C=\{i\colon\overline{x}_{i}\neq x^{0}_{i}\}$ and define a temporary set $U^{\prime}=\{i\colon\overline{x}_{i}=1\}$. Note that it is possible that $\lvert C\rvert\neq k$. In this case, we greedily modify $C$ to ensure that $U^{\prime}$ is a valid solution of GP-kkR. We call this last step fixing $C$.
To boost the probability of success, the second and third steps of the algorithm are repeated $\mathcal{O}(1/\varepsilon\log(1/\varepsilon))$ times and we return the solution with the largest objective function value. The pseudocode of our concrete algorithms for $k$-Densify and Max-Cut-kkR are presented in Appendices C.4 and C.6, resp.
Our main results. Before analyzing the SDP-based algorithm, let us state our main results obtained through the SDP-based algorithm for DS-kkR and Max-Cut-kkR. For these problems, we provide a complete analysis of the constant factor approximation guarantees for $k=\Omega(n)$. The specific value of $k$ determines the constant factor approximation ratios, and we present them in Figure 1. Note that, due to our hardness results from Corollary 3.1, this assumption is inevitable for DS-kkR if we wish to obtain $\mathcal{O}(1)$-approximation algorithms (assuming the Small Set Expansion Hypothesis [Raghavendra and Steurer, 2010]). Moreover, the assumption $k=\Omega(n)$ is necessary for the analysis of the approximation factor, similar to previous work [Feige and Langberg, 2001; Frieze and Jerrum, 1997; Han et al., 2002].
Theorem 5.1.
Let $|U|=\Omega(n)$ and $k=\Omega(n)$ where $k\leq\min\{c|U|,n-|U|\}$ for some constant $c\in(0,1)$. There exists an SDP-based randomized algorithm that runs in time $\tilde{\mathcal{O}}(n^{3.5})$ and outputs a $\mathcal{O}(1)$-approximate solution for DS-kkR with high probability.
Theorem 5.2.
If $|U|=\Omega(n)$ and $k=\Omega(n)$, there exists an SDP-based randomized algorithm that runs in time $\tilde{\mathcal{O}}(n^{3.5})$ and returns a $\mathcal{O}(1)$-approximate solution for Max-Cut-kkR with high probability.
We note that the running time is dominated by solving the semidefinite program with $\Theta(n)$ variables and $\Theta(n)$ constraints, and the time complexity is $\tilde{\mathcal{O}}(n^{3.5})$ [Jiang et al., 2020].
5.2 Heuristic
We note that the time complexity of the SDP-based algorithms is dominated by solving the SDP. By the property of our formulation, the number of constraints is $\mathcal{O}(n)$. The time complexity of solving the SDP using the state-of-the-art method is $\mathcal{O}(n^{3.5}polylog(1/\epsilon))$ [Jiang et al., 2020], where $\epsilon$ is an accuracy parameter.
As this time complexity is undesirable in practice, we also provide an efficient heuristic for GP-kkR. Our heuristic computes a set $C$, by starting with $C=\emptyset$ and greedily adding $k$ vertices to $C$. More concretely, while $|C|<k$, we greedily add vertices to $C$ by computing the value of the objective function $\mathcal{G}_{u}=\mathcal{G}\!\left(U\triangle(C\cup\{u\}\right)$, for each vertex $u\in V\setminus C$, and adding to $C$ the vertex $u^{*}$ that increases the objective function the most. The heuristic terminates when $|C|=k$.
6 Sum-of-squares algorithm
In this section, we present an optimal approximation algorithm for Max-Cut-kkR under some regimes of $k=\Omega(n)$, assuming the Unique Games Conjecture [Khot, 2002]. Our main result is stated in Theorem 6.1. Due to Corollary 4.1, the approximation algorithm from our theorem is optimal for $\tau\in(0,0.365)\cup(0.635,1)$, assuming the Unique Games Conjecture [Khot, 2002].
Theorem 6.1.
Let $\tau\in(0,1)$ be a constant, and consider Max-Cut-kkR with $k=\tau\cdot n$ refinements. Then for every $\varepsilon>0$, there is a randomized polynomial-time algorithm that runs in time $n^{O(1/poly(\varepsilon,\tau))}$, and with high probability, outputs a solution that approximates the optimal solution within a factor of $\alpha_{*}-\varepsilon$, where the value of $\alpha_{*}$ is approximately $0.858$.
Our algorithm relies on the Lasserre hierarchy formulation [Lasserre, 2002] of the problem along with the algorithmic ideas introduced by Raghavendra and Tan [2012] for solving Max-Cut-CC. We show that their method can be applied in our setting when an initial partition is given. In addition, we show that their approximation ratio still applies in our setting. Due to the complexity of the entire algorithm, here we only give an overview of our main ideas; we present the detailed algorithm and its analysis in Appendix D.
Let $\overline{\bm{x}}$ be the partition obtained by applying the randomized algorithm introduced by Raghavendra and Tan [2012]. Specifically, it is obtained by the rounding scheme on an uncorrelated Lasserre SDP solution, which we illustrate in the appendix. Let $\overline{\mathbf{SoS}}$ be the cut value of the partition induced by $\overline{\bm{x}}$, and $\operatorname{\bf Opt}$ be the optimal solution of Max-Cut-kkR. We state the main result that we obtain from applying the method of Raghavendra and Tan [2012] in Lemma 6.2. We note that the specific algorithm and its analysis are an extension of this method, with minor (though not trivial) changes.
Lemma 6.2.
Let $\tau\in(0,1)$ be a constant, and $k=\tau\cdot n$. Let $\delta>0$ be a small constant. There is a randomized polynomial-time algorithm that runs in time $n^{O(1/poly(\delta,\tau))}$, and with high probability returns a solution $\overline{\bm{x}}$ such that $\overline{\mathbf{SoS}}\geq 0.858\cdot\operatorname{\bf Opt}(1-\delta)^{2}$ and such that $\left|\sum_{i\in V}x_{i}^{0}\overline{x}_{i}-(n-2k)\right|\leq\delta^{1/48}|V|$.
We note that compared to our analysis from Section 5, the main differences are as follows: The SDP-relaxation in Section 5 also obtains a solution with objective function value approximately $0.858\cdot\operatorname{\bf Opt}$ after the hyperplane rounding. However, after the hyperplane rounding, we might have that $\lvert C\rvert$ is very far away from the desired size of $\lvert C\rvert=k$. This makes the greedy fixing step (Step 3 in the algorithm) expensive and causes a relatively larger loss in the approximation ratio. This is where the strength of the Lasserre hierarchy comes in: Lemma 6.2 guarantees that $\left|\sum_{i\in V}x_{i}^{0}\overline{x}_{i}-(n-2k)\right|\leq\delta^{1/48}|V|$, which implies that $\lvert C\rvert$ is very close to the desired size of $\lvert C\rvert=k$. Thus, we have very little loss in the approximation ratio when greedily ensuring the size of $C$ and obtain a better result overall.
Next, we use Lemma 6.2 to directly prove our main theorem.
Proof of Theorem 6.1.
Observe that $\overline{\bm{x}}$ does not necessarily satisfy the constraint on the number of refinements. We aim to obtain a partition $\bm{x}$ such that $\sum_{i\in V}x_{i}^{0}x_{i}=n-2k$.
[…]
$|s|\leq\delta^{1/48}|V|.$
Recall that $\overline{\mathbf{SoS}}$ denotes the cut value induced by $\overline{\bm{x}}$. For $i\in V$ let us denote with $\zeta(i)$ the value that node $i$ contributes to the cut. We have that
$\sum_{i\in C}\zeta(i)\leq\sum_{i\in V}\zeta(i)\leq 2\cdot\overline{\mathbf{SoS}},$ (4)
and also
$\sum_{i\in\overline{C}}\zeta(i)\leq\sum_{i\in V}\zeta(i)\leq 2\cdot\overline{\mathbf{SoS}}.$ (5)
In case we need to move $|s|$ vertices from $C$ to $\overline{C}$, we move vertices $i\in C$ with the smallest values of $\zeta(i)$. Let us use $\left\{x_{i}\right\}_{i\in V}$ to denote the final assignment obtained by moving $\frac{s}{2}$ vertices from $\overline{C}$ to $C$, and let us use $\mathbf{SoS}$ to denote the value of the cut under this assignment. In this case due to (4) we have that
$\mathbf{SoS}\geq\overline{\mathbf{SoS}}\cdot(1-|s|/|C|),$
Similarly, if we need to move $|s|$ vertices from $\overline{C}$ to $C$, we move vertices $i\in\overline{C}$ with the smallest values of $\zeta(i)$. In this case due to (5) we have that
$\mathbf{SoS}\geq\overline{\mathbf{SoS}}\cdot(1-|s|/|\overline{C}|),$
By the previous discussion, we have that
$\mathbf{SoS}\geq\overline{\mathbf{SoS}}\cdot\left(1-|s|/\min(|C|,|\overline{C}|)\right).$
Now, since $k=\tau n$ for constant $\tau\in(0,1)$, we can choose $\delta$ sufficiently small so that $\lvert V\rvert-(2\delta^{1/96}+\delta^{1/48})\lvert V\rvert\geq k\geq(2\delta^{1/96}+\delta^{1/48})\lvert V\rvert$. For such chosen $\delta$ we have
$\mathbf{SoS}\geq\overline{\mathbf{SoS}}\cdot\left(1-\frac{\delta^{1/48}n}{2\delta^{1/96}n+\delta^{1/48}n-\delta^{1/48}n}\right)\geq\overline{\mathbf{SoS}}\cdot(1-\delta^{1/96}).$
Since $\overline{\mathbf{SoS}}\geq 0.858\cdot\operatorname{\bf Opt}(1-\delta)^{2}$, we have that
$\mathbf{SoS}\geq(1-\delta^{1/96})\cdot(1-\delta)^{2}\cdot 0.858\cdot\operatorname{\bf Opt}.$
This shows that we have an $(1-3\delta^{1/96})\cdot 0.858$-approximation algorithm. By setting $\delta=(\varepsilon/3)^{96}$, we indeed obtain our desired $(1-\varepsilon)0.858$-approximation algorithm. ∎
7 Experimental evaluation
We evaluate our approximation algorithms and heuristics on a collection of real-world and synthetic datasets. Our implementation is publicly available. [^4] Our implementation is accessible through https://github.com/SijingTu/2023-OPTRefinement-code/ In this section, we only report our findings for DS-kkR, while the results for Max-Cut-kkR are in Appendix F. We aim to answer the following questions:
RQ1:
Do our algorithms increase the density of the ground-truth subgraph?
RQ2:
Do our algorithms restore a good solution after removing some nodes from the ground-truth subgraph?
RQ3:
What are the differences between the nodes identified by different algorithms?
RQ4:
Are our algorithms scalable?
RQ5:
Which algorithm has the best performance?
Datasets. We use three types of datasets: (1) Wikipedia politician page networks where the nodes represent politicians, who are labeled with their political parties, and the edges represent whether one politician appears in the Wikipedia page of another politician. (2) Networks labeled with ground-truth communities collected from SNAP [Leskovec and Krevl, 2014]. (3) Synthetic networks generated using the stochastic block model (SBM).
As for the ground-truth subgraphs, let $G[U]$ be the ground-truth subgraph induced by the set of vertices $U$; we select $U$ in different ways depending on the datasets. For the real-world datasets, we select $U$ according to the nodes’ labels: specifically, for Wikipedia networks, we select all the nodes from one political party as $U$, and for SNAP networks, we select the largest community as $U$. As for the synthetic networks generated by SBM, we always set one of the planted communities as $U$.
The statistics of the datasets and the properties of the ground-truth subgraphs are presented in both Table 2 and Table 3. More detailed descriptions, including how we set the parameters to generate our random graphs, can be found in Appendix E.1.
Table 2: Network statistics and average relative increase in density with respect to the ground-truth subgraphs. Here, $n$ and $m$ represent the number of nodes and edges in the graph; $n_{0}$ and $\rho_{0}$ denote the number of nodes and density of the ground-truth subgraph; $n^{*}$ and $\rho^{*}$ indicate the number of nodes and density of the densest subgraph. An algorithm not terminating within 2 hours is denoted by _ ; for SQD, it also indicates that the algorithm does not output a result, as no $\sigma$-quasi-elimination order exists. We set $k=10\%\,n_{0}$.
dataset Network Statistics Algorithms Blackbox Methods Baselines
$n$ $m$ $n_{0}$ $\rho_{0}$ $n^{*}$ $\rho^{*}$ Greedy SDP SQD B: Peel B: SDP Rnd
SBM-Balanced 1 000 74 940 250 37.39 1 000 74.94 0.007 0.004 0.003 0.003 0.003 -0.043
SBM-DenseSubg 1 000 81 531 250 99.68 250 99.68 -0.057 -0.057 _ -0.057 -0.057 -0.078
SBM-SparseSubg 1 000 81 192 250 24.93 250 100.10 0.059 0.065 0.036 0.065 0.065 -0.021
Wiki-ES 205 372 128 1.56 22 3.09 0.202 0.193 0.170 0.193 0.193 -0.047
Wiki-DE 768 3 059 445 3.28 129 7.80 0.270 0.277 0.254 0.277 0.277 -0.060
Wiki-GB 2 168 18 617 1 160 6.55 234 17.85 0.321 0.321 0.290 0.321 0.321 -0.072
Wiki-US 3 912 18 359 2 014 2.97 303 11.37 0.595 0.595 0.567 0.594 0.595 -0.049
SNAP-DBLP 317 080 1 049 866 7 556 2.40 115 56.57 0.436 _ 0.849 1.151 _ -0.084
SNAP-Amazon 334 863 925 872 328 2.07 97 4.80 0.075 _ _ 0.079 _ -0.089
SNAP-Youtube 1 134 890 2 987 624 2 217 4.09 1 959 45.60 1.586 _ _ 1.586 _ -0.088
Table 3: Network statistics and average relative increase in density with respect to the ground-truth subgraph with $k=10\%\,n_{0}$ random nodes moved out. Here, $n$ and $m$ represent the number of nodes and edges in the graph; $n_{0}$ and $\rho_{0}$ denote the number of nodes and density of the ground-truth subgraph; $n^{*}$ and $\rho^{*}$ indicate the number of nodes and density of the densest subgraph. An underscore (_) denotes that an algorithm does not finish within 2 hours; for SQD, it also indicates that the algorithm does not output a result, as no $\sigma$-quasi-elimination order exists. We set $\sigma=5$. Init denotes the solution containing the $k$ nodes removed from the ground-truth subgraph.
dataset Network Statistics Algorithms Blackbox Methods Baselines
$n$ $m$ $n_{0}$ $\rho_{0}$ $n^{*}$ $\rho^{*}$ Greedy SDP SQD B: Peel B: SDP Rnd Init
SBM-Balanced 1 000 74 940 250 37.39 1 000 74.94 0.110 0.110 _ 0.102 0.110 -0.045 0.110
SBM-DenseSubg 1 000 81 531 250 99.68 250 99.68 0.112 0.112 0.112 0.112 0.112 -0.078 0.112
SBM-SparseSubg 1 000 81 192 250 24.93 250 100.10 0.112 0.109 _ 0.082 0.111 -0.015 0.111
Wiki-ES 205 372 128 1.56 22 3.09 0.424 0.420 0.389 0.408 0.420 -0.069 0.200
Wiki-DE 768 3 059 445 3.28 129 7.80 0.378 0.380 0.358 0.379 0.380 -0.084 0.098
Wiki-GB 2 168 18 617 1 160 6.55 234 17.85 0.421 0.421 0.370 0.421 0.421 -0.085 0.084
Wiki-US 3 912 18 359 2 014 2.97 303 11.37 0.775 0.774 0.735 0.775 0.775 -0.035 0.125
SNAP-DBLP 317 080 1 049 866 7 556 2.40 115 56.57 0.542 _ 1.046 1.417 _ -0.092 0.109
SNAP-Amazon 334 863 925 872 328 2.07 97 4.80 0.160 _ _ 0.134 _ -0.097 0.098
SNAP-Youtube 1 134 890 2 987 624 2 217 4.09 1 959 45.60 1.893 _ _ 1.894 _ -0.096 0.135
Algorithms. We evaluate six algorithms for solving DS-kkR. We use SDP from Theorem 5.1, Greedy from Section 5.2, and two methods based on the black-box reduction from Theorem 3.4. In the black-box reduction, we use two solvers for DkkS: the peeling algorithm of Asahiro et al. [2000] and the SDP algorithm of Feige and Langberg [2001], which we denote as B: Peel and B: SDP, respectively. We also compare our methods with the SQD ($\sigma$-quasi-densify) algorithm [Matakos and Gionis, 2022], which solves $k$-Densify, where we set $\sigma=5$ as in the original paper. We also consider a baseline that picks $k$ random nodes (Rnd); for Rnd, we repeat the procedure $5$ times and compute the average results.
Evaluation. We evaluate all methods with respect to the relative increase of density, where we compare to the density of the initial subgraph $U$. Formally, assuming that a method $\mathcal{A}$ returns a subgraph $U_{\!\mathcal{A}}$, the relative increase of density with respect to the density of $U$ is ${(d(U_{\!\mathcal{A}})-d(U))}/{d(U)}$. We say that method $\mathcal{A}$ performs $X$ times better than method $\mathcal{B}$, if the relative increase of density from the subgraph returned by $\mathcal{A}$ is $X$ times larger than $\mathcal{B}$.
Initialization using ground-truth subgraphs. We begin by examining whether our algorithms indeed increase the density of a given subgraph. For this purpose, we initialize $U$ as a ground-truth subgraph from our dataset and apply the algorithms directly to the instance $(G,U,k)$. The relative increases in densities compared to $d(U)$ are presented in Table 2 and Figure 4.
As shown in Table 2 and Figure 4, all algorithms increase the density of the ground-truth subgraph in most datasets. Specifically, in Table 4, on the datasets with larger ground-truth subgraphs, i.e., Wiki-US, SNAP-DBLPand SNAP-Youtube, the relative increase of densities surpasses $40\%$. The only outlier in which the algorithms decrease the density is SBM-DenseSubg; this is because the ground truth subgraph is already the densest subgraph. These findings indicate that our algorithms are capable of identifying denser subgraphs over the input instances, addressing RQ1.
(a) Wiki-GB (b) Wiki-US (c) SNAP-DBLP (d) SNAP-Youtube
Figure 4: Relative increase of density for varying values of $k$. We initialized $U$ as a ground-truth subgraph. No results of SDP and B: SDP are reported on SNAP-DBLPand SNAP-Youtube due to the prohibitive time complexity.
(a) Wiki-GB (b) Wiki-US (c) SNAP-DBLP (d) SNAP-Youtube
Figure 5: Relative increase of density for varying values of $k$. We initialized $U$ as a ground-truth subgraph with $k$ nodes removed uniformly at random. No results of SDP and B: SDP are reported on SNAP-DBLPand SNAP-Youtube due to the prohibitive time complexity.
Initialization via ground-truth subgraphs with $k$ nodes removed. Next, we study whether our algorithms can uncover subgraphs of densities equal to or greater than those of the ground-truth subgraphs after they were slightly perturbed. For this purpose, we initialize $\hat{U}$ as a ground-truth subgraph $U$ with $k$ random nodes removed. We then apply the algorithms to the instance $(G,\hat{U},k)$. We let Init denote the solution containing the $k$ nodes that are removed to obtain $\hat{U}$. Our results are presented in Table 3 and Figure 5, which show the average relative increase in density with respect to d(U^)d($\hat${U}) across 5 initializations. In Figure 5 we also show the corresponding standard deviations.
Table 3 and Figure 5 show that, across various datasets, all algorithms consistently find subgraphs that are denser than Init for all values of $k$. Additionally, the difference in density between the algorithms and Init increases as more nodes are removed (i.e., larger $k$). Our observation suggests that our algorithms are as good, if not better, at finding subgraphs with comparable densities, and this observation addresses RQ2.
Next, study whether our algorithms can reconstruct ground-truth subgraphs. We again start with the ground-truth subgraphs and randomly remove $k$ nodes. Then, we run our algorithms to see if they can identify the removed $k$ nodes or if they select different nodes that were not part of the ground-truth subgraph. The results are reported in Figure 6, which illustrates the Jaccard similarity among the nodes selected by the algorithms and those identified by Init. The Jaccard similarity between two sets $A$ and $B$ is defined as $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
Figure 6 shows distinct behaviors of the algorithms for both the real-world and the synthetic datasets. We start by analyzing the results on the Wikipedia datasets. We notice that these datasets contain sparse ($\rho_{0}$ is around half of $\rho^{*}$) and large (around half of the nodes) ground-truth subgraphs, and many nodes within the ground-truth subgraph maintain more connections to the nodes outside of this subgraph. This structure leads all the algorithms to select different nodes than Init. Moreover, the nodes selected by SDP, B: SDP, and B: Peel have a high degree of similarity in all three datasets, whereas Greedy demonstrates a distinct pattern in Wiki-DE. This distinction can be attributed to the “peeling” routine inherent in the prior three algorithms, which iteratively removes nodes with the lowest degree starting from the graph from the intermediate step. On the other hand, Greedy iteratively adds nodes with the highest degrees. The different procedures thus result in very different node selections in some datasets.
Next, we analyze the results on the synthetic datasets. Recall that in SBM-SparseSubg, despite the ground-truth subgraph being sparse, each node within this subgraph maintains more connections internally than with nodes outside (in expectation). Under this construction, Greedy (which does not contain the “peeling” routine) and B: Peel (which highly relies on the “peeling” routine) exhibit significant performance differences. Greedy exhibits high similarity to Init, indicating that it finds the same set of $k$ removed nodes. However, B: Peel finds a different set of nodes compared to Init. Combined with its superior performance over Greedy on the SNAP-DBLP dataset in Figure 4 and Figure 5, this behavior suggests that B: Peel tends to find globally denser structures when the ground-truth subgraph is much sparser compared to the densest subgraph. On the other hand, Greedy effectively uncovers locally dense structures, albeit at the expense of overlooking globally denser structures that are less interconnected within the local subgraph.
We now look into SBM-DenseSubg. All algorithms show high similarity to Init, which can be explained as the ground-truth subgraph is the densest subgraph and is much denser than other subgraphs. Hence, any sufficiently good algorithm should recover this ground-truth subgraph. Lastly, we look into SBM-Balanced, where only B: Peel shows a different selection compared to Init, further implying its “non-local” property.
We remark that the performance of SDP and B: SDP is between B: Peel and Greedy. Specifically, they find the removed nodes for SBM-Balanced, and find nodes that belong to a denser subgraph for SBM-SparseSubg. As SQD’s choice is generally different from all the other algorithms, we present a more detailed analysis of SQD in Appendix E.4.
As anticipated, the algorithms exhibit varied node selection behaviors across different datasets. Greedy favors nodes with stronger connections to the initial subgraph. In contrast, B: Peel is inclined to identify nodes with globally higher connectivity. B: SDP and SDP offer a balance between these frameworks, addressing RQ3.
Refer to caption Refer to caption Refer to caption
(a) Wiki-DE (b) Wiki-GB (c) Wiki-US
Refer to caption Refer to caption Refer to caption
(d) SBM-Balanced (e) SBM-DenseSubg (f) SBM-SparseSubg
Figure 6: Jaccard similarities between the nodes selected by different algorithms. We initialized $U$ as a ground-truth subgraph with $k$ nodes removed uniformly at random. Init denotes the solution containing the $k$ nodes removed from the ground-truth subgraph. The similarity matrix is symmetric along the diagonal; to make a clean presentation, we only show its upper triangle. Here we set $k=50$.
Scalability of the algorithms. In Figure 7, we present the running times of our algorithms for the four largest datasets. We observe that SDP and B: SDP are not scalable due to the high cost of solving the SDP. SQD also becomes prohibitive for larger datasets, exhibiting super-linear running times as $k$ increases. The baseline approach Greedy is also not scalable on the largest dataset, SNAP-Youtube, taking around $20$ minutes. The most scalable algorithm is B: Peel, which runs in $100$ seconds on SNAP-Youtube with one million nodes. This is because the core procedure of B: Peel only requires repeatedly removing nodes with the lowest degree, and the entire algorithm can be implemented in time $O(n+m)$. In conclusion, we find that B: Peel is the most scalable algorithm among all algorithms, addressing RQ4.
(a) Wiki-US (b) SNAP-DBLP (c) SNAP-Amazon (d) SNAP-Youtube
Figure 7: Running time for varying values of $k$. We initialize $U$ as a ground-truth subgraph and then remove $k$ nodes uniformly at random. We set $k=2\%,4\%,6\%,8\%,10\%n_{0}$.
Selecting the best algorithm. Last, we summarize the performance of all algorithms and select the best algorithm according to different applications.
For smaller datasets, such as Wikipedia and SBM datasets, the differences among the algorithms are marginal. However, SDP, B: SDP, and Greedy show a slight advantage over B: Peel, as shown in Table 2 and Table 3.
For mid-sized datasets such as SNAP-DBLP and SNAP-Amazon, SDP and B: SDP fail to produce results due to the computational complexity of solving a semidefinite program. In terms of output quality, B: Peel is only outperformed by Greedy on SNAP-Amazon, as shown in Table 3. However, on SNAP-DBLP, B: Peel’s performance is 2.5 times better than that of Greedy; on SNAP-Youtube, B: Peel also marginally outperforms Greedy. This is because B: Peel is more efficient at identifying globally denser subgraphs.
To conclude, B: Peel emerges as the preferable choice for tasks with strict scalability demands. In scenarios involving smaller graphs, SDP, B: SDP and Greedy output high-quality results. For all other cases, B: Peel is recommended for its balance of output quality and scalability, thus addressing RQ5.
Summary of the appendix. More details on the running times are provided in Appendices E.2.3 and E.2.4. Additional analysis in Appendix E.3 explores the nodes selected by different algorithms on SBM-SparseSubg. Figure 10 and Figure 11 illustrate the quality of various algorithms across a broader range of $k$ values using synthetic datasets. Last, a comprehensive analysis of SQD is presented in Appendix E.4, where SQD’s distinct performances in selecting node sets and its consistent underperformance relative to other algorithms are discussed.
8 Conclusion
In this paper, we introduce the OptiRefine framework, which characterizes problems aimed at increasing the quality of an existing solution through a small number of refinements. Our contributions are as follows: First, we generalize maximum graph-partitioning problems (MAX-GP), introducing the graph partitioning with $k$ refinements (GP-kkR) problem. Second, we develop SDP-based approaches for solving GP-kkR, providing detailed analyses of maximum cut with $k$ refinements (Max-Cut-kkR) and densest subgraph with $k$ refinements (DS-kkR). Third, we propose scalable algorithms based on a black-box reduction by leveraging the connections between our problems and DkkS and Max-Cut.
From a theoretical perspective, we identify two areas for future investigation. The first concerns whether we can apply the SDP-based approach to problems with more general parameters $c_{0}$, $c_{1}$, $c_{2}$, and $c_{3}$, which characterize problems beyond MAX-GP. The second area involves analyzing the computational complexity of Max-Cut-kkR when $k\in o(n)$, particularly given that the linear relaxation approach with pipage rounding for Max-Cut-CC achieves a $0.5$-approximation regardless of $k$’s regime [Ageev and Sviridenko, 1999].
From a practical perspective, we only briefly discussed strategies for selecting the parameter $k$, and mostly treated $k$ as a given input. We consider exploring principled methods for determining $k$ to be a worthwhile direction for future research.
Acknowledgment
We sincerely thank Per Austrin for the discussions and for providing valuable feedback on the sum-of-squares algorithm. We appreciate Antonis Matakos for sharing the source code for the SQD algorithm. This research is supported by the ERC Advanced Grant REBOUND (834862), the EC H2020 RIA project SoBigData++ (871042), the Vienna Science and Technology Fund (WWTF) [Grant ID: 10.47379/VRG23013], the Approximability and Proof Complexity project funded by the Knut and Alice Wallenberg Foundation and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Omitted content of Section 3
A.1 Proof of Lemma 3.1
Let $G=(V,E)$ and $k$ be the input for the DkkS problem. Suppose that we have an algorithm $\mathcal{A}$ for the DS-kkR problem. We apply $\mathcal{A}$ on $G$ with $U=\emptyset$ and with the same parameter $k$. The output of $\mathcal{A}$ is a set $C$ of size $|C|=k$, which we use as the solution $U$ for the DkkS problem. Since we did not make any changes to $G$, the running time of $\mathcal{A}$ on $G$ is $\mathcal{O}(T(m,n,k))$. We obtain our approximation guarantees by observing that both problems have feasible solutions consisting of exactly $k$ vertices, and since $U=\emptyset$, the optimal solutions for both problems coincide.
A.2 Proof of Corollary 3.1
We conclude from Lemma 3.1 that if there is no $\beta_{k}$-approximation algorithm for the DkkS problem, then there also does not exist a $\beta_{k}$-approximation algorithm for the DS-kkR problem with the same parameter $k$.
Lee and Ghoshal [2022] showed that the DkkS problem for $k=\tau\cdot n$ is difficult to approximate within a factor of $O(\tau\log(1/\tau))$ under the Small Set Expansion Conjecture. This implies our corollary.
A.3 Proof of Lemma 3.2
Let $C_{1}^{*}$ be the optimal set of vertices selected for the $k$-Densify problem. Let $C_{1}$ be the set of vertices chosen by an $\alpha$-approximation algorithm for the $k$-Densify problem. Observe that $|C_{1}^{*}|=|C_{1}|=k$ and ${w(E[U\cup C_{1}])}\geq\alpha{w(E[U\cup C_{1}^{*}])}$.
Let $C^{*}$ be the optimal set of vertices selected for the DS-kkR problem. Notice that $\lvert C^{*}\rvert=k$. Using the condition $k\leq c\lvert U\rvert$, we find that $|U\triangle C^{*}|\geq(1-c)|U|$ and $|U|\geq\frac{1}{(1+c)}\lvert U\triangle C^{*}\rvert$. Consequently, $|U\triangle C^{*}|\geq\frac{(1-c)}{(1+c)}\lvert U\triangle C_{1}^{*}\rvert=\frac{(1-c)}{(1+c)}\lvert U\cup C_{1}^{*}\rvert$. Notice that the equality holds because $k\leq n-\lvert U\rvert$, and hence the optimal set of vertices $C_{1}^{*}$ for $k$-Densify does not intersected with $U$, implying that $U\triangle C_{1}^{*}=U\cup C_{1}^{*}$.
Since $C_{1}^{*}$ is the optimal set of vertices for the $k$-Densify problem, we have ${w(E[U\cup C_{1}^{*}])}\geq{w(E[U\triangle C^{*}])}$.
Based on the above analysis, we can derive the following inequality that illustrates the performance of the $\alpha$-approximation algorithm for the $k$-Densify problem,
$\frac{{w(E[U\triangle C^{*}])}}{\lvert U\triangle C^{*}\rvert}$ $\leq\frac{(1+c)}{(1-c)}\cdot\frac{{w(E[U\cup C_{1}^{*}])}}{\lvert U\cup C_{1}^{*}\rvert}$ (6)
$\leq\frac{1}{\alpha}\cdot\frac{(1+c)}{(1-c)}\cdot\frac{{w(E[U\cup C_{1}])}}{\lvert U\cup C_{1}\rvert}$.
The second inequality holds since $C_{1}$ is an $\alpha$-approximation solution for the $k$-Densify problem. This concludes the proof, demonstrating that an $\alpha$-approximate solution for the $k$-Densify problem implies a $\frac{1-c}{1+c}\alpha$-approximate solution for the DS-kkR problem.
A.4 Pseudocode for solving $k$-Densify and DS-kkR with a DkkS solver (in Section 3.2)
The pseudocode of our black-box method on $k$-Densify to DkkS from Section 3.2, is illustrated as Algorithm 1.
Algorithm 1 Solve $k$-Densify using a solver for DkkS
1:$(G=(V,E,w),U,k)$, access to a solver for D(k+1)(k\!+\!1)S
2:Create a new graph $G_{1}=((V\setminus U)\cup\{i^{*}\},E_{1},w_{1})$, where $i^{*}$ corresponds to the contracted vertices in $U$
3:Set $E_{1}\leftarrow E[V\setminus U]\cup\{(i^{*},j)\colon(i,j)\in E\text{ for }i\in U,j\in V\setminus U\}$
4:Set $w_{1}(i,j)\leftarrow w(i,j)$ for alledges $(i,j)\in E[V\setminus U]$
5:for all $(i^{*},j)\in E_{1}$ do
6:  $w_{1}(i^{*},j)\leftarrow\sum_{i\in U\colon(i,j)\in E}w(i,j)$
7:end for
8:Solve D(k+1)(k\!+\!1)S in $G_{1}$ to obtain a set of vertices $U^{\prime}$ of size $k+1$
9:if $i^{*}\in U^{\prime}$ then
10:  $U^{\prime}\leftarrow U^{\prime}\setminus\{i^{*}\}$
11:else
12:  $j\leftarrow\operatorname*{arg\,min}_{j\in U^{\prime}}\text{weighted degree of }j\text{ in }G[U\cup(U^{\prime}\setminus\{i^{*}\})]$
13:  $U^{\prime}\leftarrow U^{\prime}\setminus\{j\}$
14:end if
15:return $U^{\prime}$
A.5 Proof of Lemma 3.3
We use the reduction from $k$-Densify to D(k+1)(k\!+\!1)S that we described in Section 3.1. We present the pseudocode of the reduction in Algorithm 1.
We start by defining notation. Let $G=(V,E,w)$ be the input graph for $k$-Densify. We let $G_{1}=(V_{1},E_{1},w_{1})$ be the graph defined in the reduction. Formally, we set $V_{1}=(V\setminus U)\cup\{i^{*}\}$, where $i^{*}$ is the special vertex we newly added to the graph, which represents the contracted vertices in $U$, and $E_{1}=E[V\setminus U]\cup\{(i^{*},j)\colon(i,j)\in E\mbox{ for }i\in U,j\in V\setminus U\}$. We set $w_{1}(i^{*},j)=\sum_{i\in U\colon j\in V\setminus U,(i,j)\in E}w(i,j)$.
Now let $C_{2}^{*}$ be the optimal set of vertices selected for D(k+1)(k\!+\!1)S on $G_{1}$. Let $C_{2}$ be the set of vertices selected for D(k+1)(k\!+\!1)S on $G_{1}$ by an $\alpha_{k+1}$-approximation algorithm. Notice that $|C_{2}^{*}|=|C_{2}|=k+1$. Furthermore, we let $C_{1}^{*}$ denote the optimal set of vertices selected for $k$-Densify in $G$.
By definition $\frac{{w(E_{1}[C_{2}])}}{k+1}\geq\alpha_{k+1}\frac{{w(E_{1}[C_{2}^{*}])}}{k+1}\geq\alpha_{k+1}\frac{{w(E_{1}[C_{1}^{*}\cup\{i^{*}\}])}}{k+1}$. The first inequality follows from the $\alpha_{k+1}$-approximation algorithm, the second inequality follows from ${w(E_{1}[C_{2}^{*}])}\geq{w(E_{1}[C_{1}^{*}\cup\{i^{*}\}])}$ since $C_{2}^{*}$ is the optimal solution for D(k+1)(k\!+\!1)S in $G_{1}$ and since $C_{1}^{*}\cup\{i^{*}\}$ is a feasible solution for D(k+1)(k\!+\!1)S in $G_{1}$. It follows that ${w(E_{1}[C_{2}])}\geq\alpha_{k+1}{w(E_{1}[C_{1}^{*}\cup\{i^{*}\}])}$.
Notice that if $i^{*}\in C_{2}$, then $\lvert C_{2}\setminus\{i^{*}\}\rvert=k$ and $C_{2}\setminus\{i^{*}\}$ is a feasible solution for $k$-Densify in $G$. If $i^{*}\notin C_{2}$, we have to pick a node $v\in C_{2}$ so that $C_{2}\setminus\{v\}$ becomes a feasible solution for $k$-Densify. To obtain our approximation ratios, we we make the following case distinctions.
Case 1: $i^{*}\in C_{2}$. The induced edge weights in the graph $G$ after adding $C_{2}\setminus\{i^{*}\}$ to $U$ are given by $w(E[U\cup(C_{2}\setminus\{i^{*}\})])$}. In addition, for the sum of edge weights induced by $C_{2}$ on $G_{1}$ we have that $|E_{1}[C_{2}]|={w(E[C_{2}\setminus\{i^{*}\}])}+{w(E[U,C_{2}\setminus\{i^{*}\}])}$. We then derive the following bound,
${w(E[U\cup(C_{2}\setminus\{i^{*}\})])}$ $={w(E[U])}+{w(E[C_{2}\setminus\{i^{*}\}])}+{w(E[U,C_{2}\setminus\{i^{*}\}])}$ (7)
$={w(E[U])}+|E_{1}[C_{2}]|$
$\geq{w(E[U])}+\alpha_{k+1}|E_{1}[C_{1}^{*}\cup\{i^{*}\}]|$
$\geq\alpha_{k+1}({w(E[U])}+|E_{1}[C_{1}^{*}\cup\{i^{*}\}]|)$
$=\alpha_{k+1}{w(E[U\cup C_{1}^{*}])}.$
Case 2: $i^{*}\notin C_{2}$. Let $a$ be a node in $C_{2}$ with smallest weighted degree in $G_{1}[C_{2}]$, i.e., $a=\arg\min_{u\in C_{2}}\sum_{v\in C_{2}}w_{1}(u,v)$. The weight of induced edges after adding $C_{2}\setminus\{a\}$ to $U$, i.e., $w(E[U\cup(C_{2}\setminus\{a\})])$}, is then at least ${w(E[U])}+\frac{k-1}{k+1}{w(E_{1}[C_{2}\cup\{i^{*}\})}$. Now we obtain that:
${w(E[U\cup(C_{2}\setminus\{a\})])}$ $={w(E[U])}+{w(E[C_{2}\setminus\{a\}])}+{w(E[U,C_{2}\setminus\{a\}])}$ (8)
$={w(E[U])}+|E_{1}[(C_{2}\setminus\{a\})\cup\{i^{*}\}]|$
$\geq{w(E[U])}+{w(E_{1}[C_{2}\setminus\{a\}])}$
$\geq{w(E[U])}+\frac{k-1}{k+1}{w(E_{1}[C_{2}])}$
$\geq{w(E[U])}+\frac{k-1}{k+1}\alpha_{k+1}{w(E_{1}[C_{2}^{*}])}$
$={w(E[U])}+\frac{k-1}{k+1}\alpha_{k+1}{w(E_{1}[C_{1}^{*}\cup\{i^{*}\}])}$
$\geq\frac{k-1}{k+1}\alpha_{k+1}({w(E[U])}+{w(E_{1}[C_{1}^{*}\cup\{i^{*}\}])}).$
$=\frac{k-1}{k+1}\alpha_{k+1}{w(E[U\cup C_{1}^{*}])}.$
As $w(E[U\cup C_{1}^{*}])$} is the optimal objective function value for $k$-Densify, we have finished the proof.
A.6 Proof of Theorem 3.4
The theorem follows directly from combining Lemmas 3.3 and 3.2.
Appendix B Omitted content of Section 4
B.1 Proof of Lemma 4.1
Let $G=(V,E,w)$ and $k$ be an input for Max-Cut-CC. The goal is to find $C\subseteq V$ such that $\mathrm{cut}\!\left(C\right)$ is maximized. Suppose we have an algorithm $\mathcal{A}$ for Max-Cut-kkR. We apply $\mathcal{A}$ with input $(G=(V,E,w),U=\emptyset,k)$. Now $\mathcal{A}$ returns a set $C$ of size $|C|=k$ to maximize $\mathrm{cut}\!\left(U\triangle C\right)$. Given that $U=\emptyset$, $\mathrm{cut}\!\left(U\triangle C\right)=\mathrm{cut}\!\left(C\right)$.
As the constraint and the objective function for these two problems are the same, if $\mathcal{A}$ is $\alpha_{k}$-approximation algorithm for Max-Cut-kkR, then it is $\alpha_{k}$-approximation for Max-Cut-CC. As we directly use $C$ as the solution for Max-Cut-CC, the running time is $\mathcal{O}(T(m,n,k))$.
B.2 Proof of Corollary 4.1
This is a direct result from Lemma 4.1 and the hardness results of Austrin and Stanković [2019].
B.3 Pseudocode for the black-box-based algorithm for Max-Cut-kkR from Section 4.1
The pseudocode for black-box-based algorithm for Max-Cut-kkR, assuming access to a solver to Max-Cut. As we discussed in Section 4.1, is illustrated in Algorithm 2.
Algorithm 2 An algorithm for Max-Cut-kkR, assuming access to a black-box Max-Cut solver
1:$(G(V,E,w),U,k)$, access to a Max-Cut solver
2:Let $C\leftarrow\{i\colon x_{i}\neq x^{0}_{i}\}$ be the solution of applying our black-box Max-Cut solver on $G$
3:Fixing $C$: Set $C^{{}^{\prime\prime}}\leftarrow C$
4:while $\lvert C^{{}^{\prime\prime}}\rvert\neq k$ do
5:  if $\lvert C^{{}^{\prime\prime}}\rvert<k$ then
6:   $u^{*}\leftarrow\arg\max_{u\in V\setminus C^{{}^{\prime\prime}}}\mathrm{cut}\!\left(U\triangle(C^{{}^{\prime\prime}}\cup\{u\})\right)$
7:   $C^{{}^{\prime\prime}}\leftarrow C^{{}^{\prime\prime}}\cup\{u^{*}\}$
8:  else if $\lvert C^{{}^{\prime\prime}}\rvert>k$ then
9:   $u^{*}\leftarrow\arg\max_{u\in C^{{}^{\prime\prime}}}\mathrm{cut}\!\left(U\triangle(C^{{}^{\prime\prime}}\setminus\{u\})\right)$
10:   $C^{{}^{\prime\prime}}\leftarrow C^{{}^{\prime\prime}}\setminus\{u^{*}\}$
11:  end if
12:end while
13:return $C^{{}^{\prime\prime}}$
B.4 Proof of Theorem 4.2
Before we prove the theorem, we observe the following symmetry property: $\mathrm{cut}\!\left(U\triangle(V\setminus C)\right)=\mathrm{cut}\!\left(U\triangle C\right)$. We formally state it in Lemma B.1. This implies that in the approximation ratio analysis, we only need to consider the case when $k\leq n/2$. Concretely, if $k>n/2$, we simply run the algorithm with cardinality constraint equal to nkn-k. Assume that we get $C$ as the solution for Max-Cut-kkR. Then, we can return $V\setminus C$ as the desired solution with the cardinality constraint satisfied. This solution gives us the same cut value as selecting $C$ as solution. In Figure 3, we illustrate the Max-Cut-kkR problem on a graph and its symmetry property.
Lemma B.1.
$\mathrm{cut}\!\left((V\setminus U)\triangle C\right)=\mathrm{cut}\!\left(U\triangle(V\setminus C)\right)=\mathrm{cut}\!\left(U\triangle C\right)$ for any $C\subseteq V$ and $U\subseteq V$.
Proof.
For any set $A\subseteq V$, we denote $V\setminus A$ by $\overline{A}$. Our goal is to prove that $\overline{U}\triangle C=U\triangle\overline{C}=\overline{U\triangle C}$, since this condition directly implies that $\mathrm{cut}\!\left(U\triangle C\right)=\mathrm{cut}\!\left(U\triangle\overline{C}\right)=\mathrm{cut}\!\left(\overline{U}\triangle C\right)$. We expand the three formulations to show that all of them are equal.
First, let us expand $\overline{U}\triangle C$:
$\overline{U}\triangle C$ $=(\overline{U}\setminus C)\cup(C\setminus\overline{U})$ (9)
$=(\overline{U}\cap\overline{C})\cup(C\cap U)$
$=(\overline{U}\cup C)\cap(\overline{C}\cup U)$
$=(U\cup\overline{C})\cap(C\cup\overline{U})$
Second, let we expand $U\triangle\overline{C}$:
$U\triangle\overline{C}$ $=(U\setminus\overline{C})\cup(\overline{C}\setminus U)$ (10)
$=(U\cap C)\cup(\overline{C}\cap\overline{U})$
$=(U\cup\overline{C})\cap(C\cup\overline{U})$
Finally, let us expand the right-hand side:
$\overline{U\triangle C}$ $=\overline{(U\setminus C)\cup(C\setminus U)}$ (11)
$=\overline{U\setminus C}\cap\overline{C\setminus U}$
$=\overline{U\cap\overline{C}}\cap\overline{C\cap\overline{U}}$
$=(\overline{U}\cup C)\cap(\overline{C}\cup U)$ $=(U\cup\overline{C})\cap(C\cup\overline{U})$
We can see that $\overline{U}\triangle C=U\triangle\overline{C}=\overline{U\triangle C}$. ∎
Let $(G=(V,E,w),U,k=\tau n)$ be an instance of Max-Cut-kkR, and let $(U,\overline{U})$ be the initial cut of the graph $G$. Due to the discussion and the lemma above, we only consider the case when $k\leq n/2$, hence, $\tau\leq 1/2$ in the following analysis.
After running the Max-Cut solver on $G$ (without the given cut), we obtain a cut $(U^{\prime},\overline{U^{\prime}})$. Set $C$ to be one of the two sets $U^{\prime}\triangle U$ and $U^{\prime}\triangle\overline{U}$ such that the cardinality of $C$ is at most $\frac{n}{2}$.
Let $\mu$ be such that $\lvert C\rvert=\mu n$. Now we make changes to $C$ until we have that $\lvert C\rvert=k$. If $\lvert C\rvert>k$ this is done by greedily removing the vertices from $C$ which have the smallest contribution to the cut; if $\lvert C\rvert<k$, this is done by greedily adding the vertices to $C$ which increase the cut the most. We present the pseudocode of our reduction in Algorithm 2.
For each node $a\in V$, let $\mathrm{cut}_{a}\!\left(U\right)$ be the sum of edge weights of node $a$ in the cut $(U,\overline{U})$, i.e., $\mathrm{cut}_{a}\!\left(U\right)=\sum_{(a,v)\in E[U,\overline{U}]}w(a,v)$.
We will calculate how much the greedy algorithm reduces to the approximation ratio. To do so, we make case distinctions on $\lvert C\rvert>k$ or $\lvert C\rvert<k$, and we give the following Lemma B.2.
Lemma B.2.
Let $\tau$ be such that $k=\tau n$ and assume a Max-Cut solver returns the partition of the cut with $\lvert C\rvert=\mu n$, where $\mu\in(0,1]$. Suppose the initial cut has cut value $M_{0}$. After conducting the greedy procedure to meet the cardinality constraint, the sum of edge weights of the cut is lower bounded by $(1-\Theta(\frac{1}{n}))\min\{\frac{\tau^{2}}{\mu^{2}},\frac{(1-\tau)^{2}}{(1-\mu)^{2}}\}M_{0}$.
Proof.
Let $M$ be the sum of edge weights in the cut at any stage of the greedy procedure. Note that initially $M$ is equal to $M_{0}$, and let the initial $C$ be equal to $C_{0}$. At the $i$th round of the greedy algorithm, $M$ becomes $M_{i}$, and $C$ becomes $C_{i}$.
Case 1: $\lvert C\rvert>\tau n$. We notice that $\sum_{u\in C}\mathrm{cut}_{u}\!\left(U\triangle C\right)\leq 2M$. Since the algorithm always selects a node in $C$ to maximize the cut for the next round, i.e., $u^{*}=\arg\max_{u\in C}\mathrm{cut}\!\left(U\triangle(C\setminus\{u\})\right)$, the cut value drops at most $\frac{2M}{\lvert C\rvert}$.
It follows from the previous analysis that $M_{i}\geq M_{i-1}-\frac{2M_{i-1}}{\lvert C_{i-1}\rvert}$. Thus at the $t$-th step,
[…]
Note that $\lvert C_{i-2}\rvert=\lvert C_{i}\rvert+2$, and in the end $|C_{t}|=\tau n$. Therefore, we can cancel out most of the terms and obtain:
$M_{t}$ $\geq\frac{(\lvert C_{t-2}\rvert-2)(\lvert C_{t-1}\rvert-2)}{\lvert C_{0}\rvert\lvert C_{1}\rvert}M_{0}$
$=\frac{\tau n(\tau n-1)}{\mu n(\mu n-1)}M_{0}$
$\geq\frac{\tau^{2}-\frac{\tau}{n}}{\mu^{2}}M_{0}$
$=\frac{\tau^{2}}{\mu^{2}}\left(1-\frac{1}{n\tau}\right)M_{0}.$
Case 2: $\lvert C\rvert<\tau n$. Similarly, we notice that $\sum_{u\in V\setminus C}\mathrm{cut}_{u}\!\left(U\triangle C\right)\leq 2M$. As the algorithm always selects a node in $V\setminus C$ to maximize the cut for the next round, i.e., $u^{*}=\arg\max_{u\in V\setminus C}\mathrm{cut}\!\left(U\triangle(C\cup\{u\})\right)$, the cut value drops at most $\frac{2M}{n-\lvert C\rvert}$.
It follows from the previous analysis that $M_{i}\geq M_{i-1}-\frac{2M_{i-1}}{n-\lvert C_{i-1}\rvert}$. Thus at the $t$-th step,
[…]
Note that $\lvert C_{i-2}\rvert=\lvert C_{i}\rvert-2$, and in the end $\lvert C_{t}\rvert=\tau n$. Thus,
$M_{t}$ $\geq\frac{(n-\lvert C_{t-2}\rvert-2)(n-\lvert C_{t-1}\rvert-2)}{(n-\lvert C_{0}\rvert)(n-\lvert C_{1}\rvert)}M_{0}$
$=\frac{(n-\lvert C_{t}\rvert)(n-\lvert C_{t}\rvert-1)}{(n-\lvert C_{0}\rvert)(n-\lvert C_{0}\rvert+1)}M_{0}$
$=\frac{(n-\tau n)(n-\tau n-1)}{(n-\mu n)(n-\mu n+1)}M_{0}$
$=\frac{(1-\tau)(1-\tau-1/n)}{(1-\mu)(1-\mu+1/n)}M_{0}$
$=\frac{(1-\tau)^{2}}{(1-\mu)^{2}}\frac{(1-\mu)(1-\tau-1/n)}{(1-\tau)(1-\mu+1/n)}M_{0}$
$=\frac{(1-\tau)^{2}}{(1-\mu)^{2}}\frac{(1-\mu)\cdot n-\frac{(1-\mu)}{1-\tau})}{(1-\mu)n+1}M_{0}$
$\geq\frac{(1-\tau)^{2}}{(1-\mu)^{2}}\left(1-\frac{\frac{2-\tau-\mu}{1-\tau}}{(1-\mu)n+1}\right)M_{0}.$
Combining the bounds of these two cases, we obtain that $M_{t}\geq\min\left\{\frac{\tau^{2}}{\mu^{2}}(1-\frac{1}{n\tau}),\frac{(1-\tau)^{2}}{(1-\mu)^{2}}\left(1-\frac{\frac{2-\tau-\mu}{1-\tau}}{(1-\mu)n+1}\right)\right\}M_{0}$. Both $\frac{1}{n\tau}$ and $\frac{\frac{2-\tau-\mu}{1-\tau}}{(1-\mu)n+1}$ are in $\Theta(\frac{1}{n})$. We thus conclude $M_{t}\geq(1-\Theta(\frac{1}{n}))\min\left\{\frac{\tau^{2}}{\mu^{2}},\frac{(1-\tau)^{2}}{(1-\mu)^{2}}\right\}M_{0}$. ∎
Lemma B.2 implies that, if the Max-Cut solver obtains any constant approximation ratio result, and that $\tau$ is a constant (i.e., $k=\Omega(n)$). We obtain a constant approximation algorithm for Max-Cut-kkR. Specifically, let the approximation ratio for Max-Cut be $\alpha$ (for instance, $\alpha>0.87$ for the algorithm of Goemans and Williamson [1994]), the approximation for Max-Cut-kkR becomes at least $(1-\Theta(\frac{1}{n}))\min\{4\tau^{2},(1-\tau)^{2}\}\alpha$ in the worst case, where we use that $\mu^{2}\leq\frac{1}{4}$ and $(1-\mu)^{2}\leq 1$. The inequality $\mu^{2}\leq\frac{1}{4}$ comes from how we pick $C$: we always pick $C$ with the smaller cardinality. The approximation ratio in the theorem follows since $k=\tau n$. The time complexity comes from solving the Max-Cut problem using a blackbox solver and constructing and maintaining a priority queue for the greedy procedure.
Appendix C Omitted content of Section 5
In this section, we provide the content omitted from Section 5. The section is structured as follows: We begin by outlining the general SDP-based algorithm framework in Sections C.1, C.2, and C.3. Following this, we give a detailed analysis for $k$-Densify and Max-Cut-kkR. Specifically, the pseudocode for DS-kkR is presented in Section C.4, and the pseudocode for Max-Cut-kkR is in Section C.6. We then present the proof of Theorem 5.1, the main theorem for DS-kkR and $k$-Densify, in Section C.5, and the proof of Theorem 5.2, the main theorem for Max-Cut-kkR, in Section C.7.
C.1 Useful tools for the analysis
Let us first introduce additional notation and lemmas that we will repeatedly use in this section. We define the constants
$\alpha=\min_{0\leq\theta\leq\pi}\frac{2}{\pi}\frac{\theta}{1-\cos\theta}>0.87856,$ (12)
and
$\beta=\min_{0\leq\theta<\arccos(-1/3)}\frac{2}{\pi}\frac{2\pi-3\theta}{1+3\cos\theta}>0.79607.$ (13)
We also introduce the binary function $\operatorname{sgn}\colon\mathbb{R}\rightarrow\{-1,1\}$ such that $\operatorname{sgn}(x)=1$ if $x\geq 0$, and $\operatorname{sgn}(x)=-1$ if $x<0$. Additionally, we will use the following lemmas.
Lemma C.1 (Lemma 2.1 in Goemans and Williamson [1994]).
It holds for any $y\in[-1,1]$ that $\arccos(y)\geq\alpha\pi\frac{1}{2}(1-y)$, where $\alpha$ is defined in Equation (12).
Lemma C.2 (Lemma 1.2 in Goemans and Williamson [1994]).
Let $\bm{v}_{i}$ and $\bm{v}_{j}$ be vectors on the unit sphere $\mathcal{S}_{n}$, let $\bm{r}$ be a unit vector that is drawn uniformly randomly from $\mathcal{S}_{n}$, then $\Pr(\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\neq\operatorname{sgn}(\bm{v}_{j}\cdot\bm{r}))=\frac{1}{\pi}\arccos(\bm{v}_{i}\cdot\bm{v}_{j})$.
Lemma C.3 (Lemma 2.3 in Goemans and Williamson [1994]).
Let $\bm{v}_{i}$, $\bm{v}_{j}$ and $\bm{v}_{k}$ be three vectors on the unit sphere $\mathcal{S}_{n}$, let $y_{1}=\bm{v}_{i}\cdot\bm{v}_{j}$, $y_{2}=\bm{v}_{j}\cdot\bm{v}_{k}$ and $y_{3}=\bm{v}_{i}\cdot\bm{v}_{k}$. It follows that $1-\frac{1}{2\pi}(\arccos(y_{1})+\arccos(y_{2})+\arccos(y_{3}))\geq\frac{\beta}{4}(1+y_{1}+y_{2}+y_{3})$, where $\beta$ is defined in Equation (13).
C.2 Analysis of the hyperplane rounding
In this subsection, we analyze the second step of the SDP-based algorithm for GP-kkR. In particular, we derive bounds that hold for the GP-kkR after hyperplane rounding. Our bounds are presented in Lemma C.4. Similar to Frieze and Jerrum [1997]; Feige and Langberg [2001]; Han et al. [2002], we introduce the random variable $Z$ that controls the number of nodes that change the position and the objective value. Then we derive a lower bound on $Z$ with high probability using Markov’s inequality. We present the proof at the end of this section.
In our proof, we use the following notation. We let $\overline{\bm{x}}\in\{-1,1\}^{n}$ be the indicator vector obtained from the hyperplane rounding; notice that $\overline{\bm{x}}$ does not necessarily obey the cardinality constraint. We set $\overline{\mathbf{SDP}}$ to the objective value of GP-kkR induced by $\overline{\bm{x}}$. Let $\operatorname{\bf Opt}$ be the optimal value of GP-kkR. We let $\lambda=\frac{\overline{\mathbf{SDP}}}{\operatorname{\bf Opt}}$ and $\varphi=\frac{\mathbb{E}\left[\overline{\mathbf{SDP}}\right]}{\operatorname{\bf Opt}}$. Notice that $\lambda$ is a random variable and $\mathbb{E}\left[\lambda\right]=\varphi$. We let $\tau$ be such that $k=\tau n$ and notice that $\tau=\Omega(1)$ since $k=\Omega(n)$. Moreover, we let $C=\{i\colon\overline{x}_{i}\neq x^{0}_{i}\}$ and $U^{\prime}=\{i\colon\overline{x}_{i}=1\}$.
Lemma C.4.
Let $\gamma$ and $\eta$ be two constants such that $\gamma\in[0.1,5]$ and $\eta\in[\frac{1-\alpha+0.01}{\alpha},+\infty)$. Let $Z$ be the random variable such that
$Z=\lambda+\gamma\eta\frac{n-\lvert C\rvert}{n-k}+\gamma\frac{\lvert C\rvert(2k-\lvert C\rvert)}{n^{2}}.$
After repeating the hyperplane rounding at least $N=\frac{1-p+\epsilon p}{\epsilon p}\log(\frac{1}{\epsilon})$ times, where $p$ is a constant lower bounded by $0.0005$, it holds that with probability at least $1-\epsilon$, $Z\geq(1-\epsilon)[\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)]$.
To prove Lemma C.4, we must introduce several bounds on $\varphi$ and the expected value of $\lvert C\rvert$. We begin with the bound on $\varphi$. For Lemma C.4 to be valid, we rely on the fact that $\varphi>0$ for the parameter settings in Table 1. The values of $\varphi$ for $k$-Densify and Max-Cut-kkR are presented in Corollaries C.1 and C.2, respectively. For the $\varphi$ values related to VC-kkR and Max-Uncut-kkR, we refer readers to [Feige and Langberg, 2001].
Lemma C.5.
Let $\tilde{\varphi}$ be the maximum value such that the inequality
$(c_{0}+c_{1}+c_{2}+c_{3})-\nicefrac{{2}}{{\pi}}\cdot(c_{1}\arccos(\bm{v}_{0}\cdot\bm{v}_{i})+c_{2}\arccos(\bm{v}_{0}\cdot\bm{v}_{j})+c_{3}\arccos(\bm{v}_{i}\cdot\bm{v}_{j}))$
$\geq\tilde{\varphi}\cdot(c_{0}+c_{1}\bm{v}_{0}\cdot\bm{v}_{i}+c_{2}\bm{v}_{0}\cdot\bm{v}_{j}+c_{3}\bm{v}_{i}\cdot\bm{v}_{j})$
holds for every $\bm{v}_{0},\bm{v}_{i},\bm{v}_{j}\in\mathcal{S}_{n}$. It holds that $\varphi\geq\tilde{\varphi}$.
Proof.
First, we observe that the condition implies that
$\sum_{i<j}w(i,j)\left((c_{0}+c_{1}+c_{2}+c_{3})-\nicefrac{{2}}{{\pi}}\cdot(c_{1}\arccos(\bm{v}_{0}\cdot\bm{v}_{i})+c_{2}\arccos(\bm{v}_{0}\cdot\bm{v}_{j})+c_{3}\arccos(\bm{v}_{i}\cdot\bm{v}_{j}))\right)$
$\geq\tilde{\varphi}\cdot\sum_{i<j}w(i,j)(c_{0}+c_{1}\bm{v}_{0}\cdot\bm{v}_{i}+c_{2}\bm{v}_{0}\cdot\bm{v}_{j}+c_{3}\bm{v}_{i}\cdot\bm{v}_{j}).$
Next, we show that
$\mathbb{E}\left[(c_{0}+c_{1}\overline{x}_{i}+c_{2}\overline{x}_{j}+c_{3}\overline{x}_{i}\overline{x}_{j})\right]$
$=(c_{0}+c_{1}+c_{2}+c_{3})-\nicefrac{{2}}{{\pi}}\cdot(c_{1}\arccos(\bm{v}_{0}\cdot\bm{v}_{i})+c_{2}\arccos(\bm{v}_{0}\cdot\bm{v}_{j})+c_{3}\arccos(\bm{v}_{i}\cdot\bm{v}_{j}))$
by using Lemma C.2 as follows,
$\mathbb{E}\left[(c_{0}+c_{1}\overline{x}_{i}+c_{2}\overline{x}_{j}+c_{3}\overline{x}_{i}\overline{x}_{j})\right]$
$=c_{0}+c_{1}\mathbb{E}\left[\operatorname{sgn}((\bm{r}\cdot\bm{v}_{0})(\bm{r}\cdot\bm{v}_{i}))\right]+c_{2}\mathbb{E}\left[\operatorname{sgn}((\bm{r}\cdot\bm{v}_{0})(\bm{r}\cdot\bm{v}_{j}))\right]+c_{3}\mathbb{E}\left[\operatorname{sgn}((\bm{r}\cdot\bm{v}_{i})(\bm{r}\cdot\bm{v}_{j}))\right]$
$=c_{0}+c_{1}(1-\nicefrac{{2}}{{\pi}}\cdot\arccos{(\bm{v}_{0}\cdot\bm{v}_{i})})+c_{2}(1-\nicefrac{{2}}{{\pi}}\cdot\arccos{(\bm{v}_{0}\cdot\bm{v}_{j})})+c_{3}(1-\nicefrac{{2}}{{\pi}}\cdot\arccos{(\bm{v}_{i}\cdot\bm{v}_{j})})$
$=c_{0}+c_{1}+c_{2}+c_{3}-\nicefrac{{2}}{{\pi}}\cdot(c_{1}\arccos{(\bm{v}_{0}\cdot\bm{v}_{i})}+c_{2}\arccos{(\bm{v}_{0}\cdot\bm{v}_{j})}+c_{3}\arccos{(\bm{v}_{i}\cdot\bm{v}_{j})}).$
Next, notice that $\operatorname{\bf Opt}\leq\sum_{i<j}w(i,j)(c_{0}+c_{1}\bm{v}_{0}\cdot\bm{v}_{i}+c_{2}\bm{v}_{0}\cdot\bm{v}_{j}+c_{3}\bm{v}_{i}\cdot\bm{v}_{j})$ and $\mathbb{E}\left[\overline{\mathbf{SDP}}\right]=\sum_{i<j}w(i,j)\mathbb{E}\left[c_{0}+c_{1}\overline{x}_{i}+c_{2}\overline{x}_{j}+c_{3}\overline{x}_{i}\overline{x}_{j}\right]$, and we get that
$\mathbb{E}\left[\overline{\mathbf{SDP}}\right]\geq\tilde{\varphi}\operatorname{\bf Opt}.$
By definition we get $\varphi\geq\tilde{\varphi}$, and we conclude the lemma.
Corollary C.1.
For $k$-Densify, it holds that $\varphi=\frac{\mathbb{E}\left[{w(E[U^{\prime}])}\right]}{\operatorname{\bf Opt}}\geq\beta$.
Proof.
First, we note that $\mathbb{E}\left[{w(E[U^{\prime}])}\right]=\mathbb{E}\left[\overline{\mathbf{SDP}}\right]$ when the problem is $k$-Densify. Next, recall that for $k$-Densify we have $c_{0}=\frac{1}{4}$, $c_{1}=\frac{1}{4}$, $c_{2}=\frac{1}{4}$, $c_{3}=\frac{1}{4}$ in Lemma C.5. Now the ratio $\beta$ is obtained by applying Lemma C.3. ∎
Corollary C.2.
For Max-Cut-kkR, it holds that $\varphi=\frac{\mathbb{E}\left[\lvert\mathrm{cut}[U^{\prime}]\rvert\right]}{\operatorname{\bf Opt}}\geq\alpha$.
Proof.
First, we note that $\mathbb{E}\left[\lvert\mathrm{cut}[U^{\prime}]\rvert\right]=\mathbb{E}\left[\overline{\mathbf{SDP}}\right]$ when the problem is Max-Cut-kkR. Next, recall that for Max-Cut-kkR we have $c_{0}=\frac{1}{2}$, $c_{1}=0$, $c_{2}=0$, $c_{3}=-\frac{1}{2}$ in Lemma C.5. Now the ratio $\alpha$ is obtained by applying Lemma C.1. ∎
[…]
Lemma C.6.
The expected size of $C$ can be bounded by $\mathbb{E}\left[\lvert C\rvert\right]\geq\alpha k$ and $\mathbb{E}\left[\lvert C\rvert\right]\leq n-n\alpha+\alpha k$.
Proof.
We start by defining a sequence of Bernoulli random variables $\{Y_{i}\}_{i=1,\ldots,n}$ such that $Y_{i}=1$ if $\overline{x}_{i}\neq x_{i}^{0}$ and, otherwise, $Y_{i}=0$. By definition, $\Pr(Y_{i}=1)=\Pr(\overline{x}_{i}\neq x_{i}^{0})$. Since $x_{i}^{0}$ is a given constant that can either be 1-1 or 11, we can write $\Pr(Y_{i}=1)$ as the summation of joint probability over $x_{i}^{0}=1$ and $x_{i}^{0}=-1$, that is
$\Pr(Y_{i}=1)$ $=\Pr(\overline{x}_{i}\neq x_{i}^{0},x_{i}^{0}=-1)+\Pr(\overline{x}_{i}\neq x_{i}^{0},x_{i}^{0}=1).$
Recall that we set $\overline{x}_{i}=1$ if $(\bm{v}_{0}\cdot\bm{r})(\bm{v}_{i}\cdot\bm{r})\geq 0$ and $\overline{x}_{i}=-1$ otherwise. Thus, $\overline{x}_{i}=\operatorname{sgn}(\bm{v}_{0}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})$, and $x_{i}^{0}\neq\overline{x}_{i}$ is equivalent to $x_{i}^{0}\neq\operatorname{sgn}(\bm{v}_{0}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})$. We make the case distinctions for $x_{i}^{0}=-1$ and $x_{i}^{0}=1$.
Case 1: $x_{i}^{0}=-1$. Then $\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})=-\operatorname{sgn}(\bm{v}_{0}x_{i}^{0}\cdot\bm{r})$, and
$\Pr(\overline{x}_{i}\neq x_{i}^{0},x_{i}^{0}=-1)$ $=\Pr(\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{0}x_{i}^{0}\cdot\bm{r})=-1,x_{i}^{0}=-1).$
Case 2: $x_{i}^{0}=1$. Then $\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})=\operatorname{sgn}(\bm{v}_{0}x_{i}^{0}\cdot\bm{r})$, and
$\Pr(\overline{x}_{i}\neq x_{i}^{0},x_{i}^{0}=1)$ $=\Pr(\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{0}x_{i}^{0}\cdot\bm{r})=-1,x_{i}^{0}=1).$
Adding up the above two formulas,
$\Pr(Y_{i}=1)$ $=\Pr(\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{0}x_{i}^{0}\cdot\bm{r})=-1)$ $\overset{(1)}{=}\frac{1}{\pi}\arccos(x_{i}^{0}\bm{v}_{0}\cdot\bm{v}_{i}),$
where $(1)$ holds according to Lemma C.2.
Note that $\lvert C\rvert=\sum_{i=1}^{n}Y_{i}$, and by the linearity of expectation, $\mathbb{E}\left[\lvert C\rvert\right]=\sum_{i=1}^{n}\mathbb{E}\left[Y_{i}\right]$. From now on, the analysis is the same as the one that appears in Feige and Langberg [2001, Lemmas 2.3 and 3.2] to obtain lower and upper bounds on $\mathbb{E}\left[\lvert C\rvert\right$]. In detail, the lower bound on $\mathbb{E}\left[\lvert C\rvert\right$] is obtained from
$\mathbb{E}\left[\lvert C\rvert\right$] $\overset{(a)}{=}\frac{1}{\pi}\sum_{i=1}^{n}\arccos(x_{i}^{0}\bm{v}_{0}\cdot\bm{v}_{i})$ $\overset{(b)}{\geq}\alpha\sum_{i=1}^{n}\frac{1-x_{i}^{0}\bm{v}_{0}\cdot\bm{v}_{i}}{2}\overset{(c)}{=}\alpha k.$ (14)
Similarly, the upper bound is obtained from $\mathbb{E}\left[\lvert C\rvert\right$]:
$\mathbb{E}\left[\lvert C\rvert\right$] $\overset{(a)}{=}\sum_{i=1}^{n}\left(1-\frac{1}{\pi}\arccos(-x_{i}^{0}\bm{v}_{0}\cdot\bm{v}_{i})\right)$ (15)
$\overset{(b)}{\leq}n-\alpha\sum_{i=1}^{n}\frac{1+x_{i}^{0}\bm{v}_{0}\cdot\bm{v}_{i}}{2}\overset{(c)}{=}(1-\alpha)n+\alpha k.$
In the above formulas, $(a)$ follows from the definition of $\lvert C\rvert$, $(b)$ follows from Lemma C.1, and $(c)$ follows from the constraint on the SDP (3).
Lemma C.7.
It holds that $\mathbb{E}\left[\lvert C\rvert(n-\lvert C\rvert)\right]\geq\alpha(n-k)k$.
Proof.
Let $R_{ij}$ denote the Bernoulli random variable indicating whether the relative locations of nodes $i$ and $j$ has changed, that is,
$R_{ij}=\begin{cases}1,&\text{if }(\overline{x}_{i}\neq\overline{x}_{j}\text{ and }x_{i}^{0}=x_{j}^{0})\text{ or }(\overline{x}_{i}=\overline{x}_{j}\text{ and }x_{i}^{0}\neq x_{j}^{0}),\\ 0,&\text{otherwise}.\end{cases}$
By definition, $\Pr(R_{ij}=1)=\Pr(\overline{x}_{i}\neq\overline{x}_{j},x_{i}^{0}=x_{j}^{0})+\Pr(\overline{x}_{i}=\overline{x}_{j},x_{i}^{0}\neq x_{j}^{0})$.
Notice that $\overline{x}_{i}=\operatorname{sgn}(\bm{v}_{0}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})$. Now $\overline{x}_{i}\neq\overline{x}_{j}$ can be equivalently written as $\overline{x}_{i}\overline{x}_{j}=-1$ and, based on the statement before, is also equivalent to $\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{j}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{0}\cdot\bm{r})^{2}=-1$. Note that the term $\operatorname{sgn}(\bm{v}_{0}\cdot\bm{r})^{2}$ is always equal to $1$ and can thus be dropped. The first term of $\Pr(R_{ij}=1)$ can thus be formulated as
$\Pr(\overline{x}_{i}\neq\overline{x}_{j},x_{i}^{0}x_{j}^{0}=1)$ $=\Pr(\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{j}\cdot\bm{r})=-1,x_{i}^{0}x_{j}^{0}=1)$
$=\Pr(\operatorname{sgn}(x_{i}^{0}\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(x_{j}^{0}\bm{v}_{j}\cdot\bm{r})=-1,x_{i}^{0}x_{j}^{0}=1).$
Similarly, $\overline{x}_{i}\overline{x}_{j}=1$ can be equivalently formulated as $\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{j}\cdot\bm{r})=1$. The second term of $\Pr(R_{ij}=1)$ can thus be formulated as
$\Pr(\overline{x}_{i}=\overline{x}_{j},x_{i}^{0}x_{j}^{0}=-1)$ $=\Pr(\operatorname{sgn}(\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\bm{v}_{j}\cdot\bm{r})=1,x_{i}^{0}x_{j}^{0}=-1)$
$=\Pr(\operatorname{sgn}(x_{i}^{0}\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(x_{j}^{0}\bm{v}_{j}\cdot\bm{r})=-1,x_{i}^{0}x_{j}^{0}=-1).$
After summing up the two terms and applying Lemma C.2 in the last equality, we get that
$\Pr(R_{ij}=1)$ $=\Pr(\overline{x}_{i}\neq\overline{x}_{j},x_{i}^{0}x_{j}^{0}=1)+\Pr(\overline{x}_{i}=\overline{x}_{j},x_{i}^{0}x_{j}^{0}=-1)$
$=\Pr(\operatorname{sgn}(\overline{x}_{i}\bm{v}_{i}\cdot\bm{r})\operatorname{sgn}(\overline{x}_{j}\bm{v}_{j}\cdot\bm{r})=-1)$
$=\frac{\arccos(x_{i}^{0}x_{j}^{0}\bm{v}_{i}\cdot\bm{v}_{j})}{\pi}.$
Recall that $C=\{i\colon\overline{x}_{i}\neq x^{0}_{i}\}$ is the set of vertices that changed sides after the hyperplane rounding; thus, $C$ is a random set and also $\lvert C\rvert$ is a random variable. Note that by definition, $\lvert C\rvert(n-\lvert C\rvert)=\sum_{i<j}R_{ij}$. Hence, by the linearity of expectation, $\mathbb{E}\left[\lvert C\rvert(n-\lvert C\rvert)\right]=\sum_{i<j}\mathbb{E}\left[R_{ij}\right]$. Now we obtain our lower bound as follows:
[…]
$\overset{(b)}{\geq}\alpha\sum_{i<j}\frac{1-x_{i}^{0}x_{j}^{0}\bm{v}_{i}\cdot\bm{v}_{j}}{2}$
$\overset{(c)}{\geq}\alpha(nk-k^{2}),$
where holds by our derivation above, $(b)$ follows by Lemma C.1, and $(c)$ holds by the constraints in the SDP (3). ∎
Proof of Lemma C.4. Now, we are ready to prove our main result for the hyperplane rounding step of our algorithm.
Proof of Lemma C.4.
We start the proof with a lower bound on $\mathbb{E}\left[Z\right]$. By applying Lemma C.6, Lemma C.7 and Lemma C.5, we get $\mathbb{E}\left[Z\right]\geq\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)$.
Next, observe that $Z<1+\gamma\eta+\gamma$, as $\varphi<1$, $\alpha<1$ and $\alpha(1-\tau)^{2}-1+2\tau<1$. By Markov’s inequality, we get $\Pr(Z\leq(1-\epsilon)(\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)))\leq\frac{1-p}{1-(1-\epsilon)p},$ for $p=\frac{\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)}{1+\gamma\eta+\gamma}$. More precisely, this inequality can obtained from
$\Pr(Z\leq(1-\epsilon)(\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)))$
$\leq\frac{1+\gamma\eta+\gamma-\mathbb{E}\left[Z\right]}{1+\gamma\eta+\gamma-(1-\epsilon)(\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau))}$
$\leq\frac{1+\gamma\eta+\gamma-[\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)]}{1+\gamma\eta+\gamma-(1-\epsilon)(\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau))}$
$\leq\frac{1-p}{1-(1-\epsilon)p}.$
Now observe that $p<1$. Additionally, we derive a lower bound on $p$, and we notice that $p>0.0005$ since
$p$ $=\frac{\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)}{1+\gamma\eta+\gamma}$
$\overset{(a)}{>}\frac{\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)}{1+\gamma\eta+\gamma}$
$\overset{(b)}{>}\frac{\gamma\eta\alpha+\gamma(\alpha-1)}{1+\gamma\eta+\gamma}$
$\overset{(c)}{\geq}\frac{0.01\gamma}{1+1.01\gamma/\alpha}$
[…]
Above, $(a)$ holds as $\varphi>0$, $(b)$ holds as $2\alpha\tau<2\tau$, $(c)$ holds as the formula obtains the minimum value by setting $\eta=\frac{1-\alpha+0.01}{\alpha}$, $(d)$ holds as the denominator is at most $2$ and the numerator is at least $0.001$ by setting $\gamma=0.1$.
As we repeat the hyperplane rounding $N$ times, and pick the largest value of $Z$ for fixed $\gamma$ and $\eta$, the probability that $Z\leq(1-\epsilon)(\beta+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau))$ is bounded from above by
$\left[\frac{1-p}{1-(1-\epsilon)p}\right]^{N}$ $\leq\left[1-\frac{1}{1+\frac{1-p}{\epsilon p}}\right]^{(1+\frac{1-p}{\epsilon p})\frac{\epsilon p}{1-p+\epsilon p}N}$
$\leq\exp\left[-\frac{\epsilon p}{1-p+\epsilon p}N\right].$
Thus if we choose $N\geq\frac{1-p+\epsilon p}{\epsilon p}\log(\frac{1}{\epsilon})$, we can guarantee the above probability is upper bounded by $\epsilon$, that is, we can guarantee $Z\geq(1-\epsilon)(\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau))$ for a small positive value $\epsilon$, with probability at least $1-\epsilon$. ∎
C.3 Analysis of fixing $C$
Next, we analyze the third step of the SDP-based algorithm, which ensures that the cardinality constraint for $C$ is met. Let the size of the selected set of nodes after the hyperplane rounding be $\mu n$, i.e., we let $\mu$ be such that $\lvert C\rvert=\mu n$. Notice that $\mu$ is a random variable and $1\geq\mu>0$. Recall that $k\in\Omega(n)$ and $k=\tau n$, where $\tau$ is an input constant. We analyze the cost of fixing $C$ to ensure that the cardinality constraint is met and illustrate how the approximation ratios of GP-kkR are computed.
Let us start from a direct implication from Lemma C.4, where we set $\epsilon=\frac{1}{\Theta(n)}$, which is small enough relative to other variables.
Corollary C.3.
Let $\operatorname{\bf Opt}$ be the optimal value of the GP-kkR problem. After repeating the hyperplane rounding $N=\Theta(n\log n)$ times, with probability at least $1-\frac{1}{\Theta(n)}$, there exists a solution $\overline{\mathbf{SDP}}$ after the hyperplane rounding such that
$\frac{\overline{\mathbf{SDP}}}{\operatorname{\bf Opt}}\geq(\varphi+\gamma\eta\alpha+\gamma(\alpha(1-\tau)^{2}-1+2\tau)-\gamma\eta(1-\mu)\frac{1}{1-\tau}-\gamma\mu(2\tau-\mu))-\frac{1}{\Theta(n)}.$ (17)
Recall that $\lambda=\frac{\overline{\mathbf{SDP}}}{\operatorname{\bf Opt}}$. Based on Corollary C.3, we can write $\lambda$ as a function of $\mu$, $\gamma$ and $\eta$ with given $\tau$, and we denote this function by $\lambda_{\tau}(\mu,\gamma,\eta)$. Then we obtain that
$\frac{\overline{\mathbf{SDP}}}{\operatorname{\bf Opt}}$ $=\lambda_{\tau}(\mu,\gamma,\eta).$ (18)
Let $\mathbf{SDP}$ be the objective value of GP-kkR after fixing $C$ through the greedy procedure. Next, we outline a general SDP-based algorithm framework for analyzing the approximation ratio of GP-kkR. For all the GP-kkR problems, we note that the ratio between $\mathbf{SDP}$ and $\overline{\mathbf{SDP}}$ is a function of $\tau=\frac{k}{n}$ and $\mu=\frac{\lvert C\rvert}{n}$. Let us introduce a function $\kappa_{\tau}(\mu)$ such that
$\frac{\mathbf{SDP}}{\overline{\mathbf{SDP}}}=\kappa_{\tau}(\mu).$ (19)
Then the approximation ratio of GP-kkR is given by
$\frac{\mathbf{SDP}}{\operatorname{\bf Opt}}=\frac{\overline{\mathbf{SDP}}}{\operatorname{\bf Opt}}\cdot\frac{\mathbf{SDP}}{\overline{\mathbf{SDP}}}=\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma,\eta),$
and the approximation ratio depends on the specific settings of $\gamma$ and $\eta$. For the worst case analysis, we can bound it by selecting the most appropriate $\gamma$ and $\eta$ by solving
$\max_{\gamma\in[0.1,5],\eta\in[\frac{1-\alpha+0.01}{\alpha},+\infty)}\min_{\mu\in[0,1]}\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma,\eta).$ (20)
C.4 Pseudocode for SDP-based algorithm for $k$-Densify and DS-kkR (in Section 5.1)
The pseudocode for the SDP-based algorithm for $k$-Densify and DS-kkR, as discussed in Section 5.1, is illustrated in Algorithm 3.
Algorithm 3 An SDP-based algorithm for $k$-Densify and DS-kkR
1:$(G(V,E,w),U,k)$, number of iterations NN
2:Step 1: Solve the SDP. Solve the SDP (3) to obtain $\bm{v}_{0},\ldots,\bm{v}_{n}$
3:for $i=1$ to NN do
4:  Step 2: Hyperplane Rounding. Obtain the indicator vector $\overline{\bm{x}}$
5:  Sample a unit vector $\bm{r}$ uniformly randomly from $\mathcal{S}_{n}$
6:  for $i=1$ to $n$ do
7:   if $(\bm{r}\cdot\bm{v}_{0})(\bm{r}\cdot\bm{v}_{i})\geq 0$ then
8:     $\overline{x}_{i}\leftarrow 1$
9:   else
10:     $\overline{x}_{i}\leftarrow-1$
11:   end if
12:  end for
13:  Set $C\leftarrow\{i\colon\overline{x}_{i}\neq x^{0}_{i}\}$ and $U^{\prime}\leftarrow\{i\colon\overline{x}_{i}=1\}$
14:  Step 3: Fixing $C$. Set $C^{{}^{\prime\prime}}\leftarrow C$
15:  if $\lvert C^{{}^{\prime\prime}}\rvert<k$ then
16:   Add $k-\lvert C^{{}^{\prime\prime}}\rvert$ arbitrary vertices from $V\setminus(C^{{}^{\prime\prime}}\cup U^{\prime})$ to $C^{{}^{\prime\prime}}$
17:  else if $\lvert C^{{}^{\prime\prime}}\rvert>k$ then
18:   if $\lvert C^{{}^{\prime\prime}}\setminus U^{\prime}\rvert\geq\lvert C^{{}^{\prime\prime}}\rvert-k$ then
19:     Remove $\lvert C^{{}^{\prime\prime}}\rvert-k$ arbitrary vertices in $C^{{}^{\prime\prime}}\setminus U^{\prime}$ from $C^{{}^{\prime\prime}}$
20:   else
21:     Remove all vertices in $C^{{}^{\prime\prime}}\setminus U^{\prime}$ from $C$
22:     Greedily choose $\lvert C^{{}^{\prime\prime}}\cap U^{\prime}\rvert-k$ minimum weighted degree nodes in subgraph $G[C^{{}^{\prime\prime}}\cap U^{\prime}]$ and them remove from $C^{{}^{\prime\prime}}$
23:   end if
24:  end if
25:end for
26:return $C^{{}^{\prime\prime}}$
C.5 Proof of Theorem 5.1
The pseudocode of our algorithm can be found in Algorithm 3. In the present section, we demonstrate that Algorithm 3 provides a constant factor approximation to the $k$-Densify problem. As implied by Lemma 3.2, this solution also offers a constant approximation for the DS-kkR problem under the theorem’s specified condition.
C.5.1 Fixing $C$
Notice that at the end of a round of hyperplane rounding, we get $C=\{i\colon x_{i}\neq x_{i}^{0}\}$, and $U^{\prime}=U\triangle C=\{i\colon x_{i}=1\}$. In addition, we use Corollary C.1 to bound $\mathbb{E}\left[{w(E[U^{\prime}])}\right]\geq\beta W_{{\small SDP}}^{*}$. Since $\lvert C\rvert$ is not necessarily equal to $k$, the algorithm needs to alter $C$. This is done by considering a new variable $C^{{}^{\prime\prime}}=C$ and making changes to $C^{{}^{\prime\prime}}$ until $\lvert C^{{}^{\prime\prime}}\rvert=k$.
To give a lower bound on $w(E[U\triangle C^{{}^{\prime\prime}}])$}, we notice that only the operation that greedily removes nodes in $C^{{}^{\prime\prime}}\cap U^{\prime}$ from $C^{{}^{\prime\prime}}$ decreases $w(E[U\triangle C^{{}^{\prime\prime}}])$}. In the worst case scenario, all $\lvert C\rvert-k$ nodes are removed. Lemma C.8 bounds the decrease of $w(E[U^{\prime}])$} after this greedy procedure.
Lemma C.8.
For $k$-Densify, if $\mu>\tau$, $\mathbf{SDP}\geq\frac{\tau^{2}}{\mu^{2}}(1-\frac{1}{\Theta(n)})\overline{\mathbf{SDP}}$, in other words, the sum of edge weights of the subgraph is lower bounded by ${w(E[U^{\prime}])}\frac{\tau^{2}}{\mu^{2}}(1-\frac{1}{\Theta(n)})$. Otherwise, $\mathbf{SDP}\geq\overline{\mathbf{SDP}}$.
Proof.
Notice that if $\tau\geq\mu$, to satisfy the cardinality constraint, $U^{\prime}$ adds more nodes, thus edges, and the sum of edge weights increases. In this case, $\mathbf{SDP}\geq\overline{\mathbf{SDP}}$.
When $\tau<\mu$, we let $U^{{}^{\prime\prime}}=U\triangle C^{{}^{\prime\prime}}$. We allow $U^{{}^{\prime\prime}}$ to change correspondingly with $C^{{}^{\prime\prime}}$. Initially, $U^{{}^{\prime\prime}}=U^{\prime}$, and we will demonstrate that, in the end, ${w(E[U^{{}^{\prime\prime}}])}\geq{w(E[U^{\prime}])}\frac{k^{2}}{\lvert C\rvert^{2}}(1-\frac{1}{k})$.
We analyze three operations that modify $C^{{}^{\prime\prime}}$. We first note that the operation which adds nodes from $V\setminus(C^{{}^{\prime\prime}}\cup U^{\prime})$ to $C^{{}^{\prime\prime}}$ only occurs when $\lvert C\rvert<k$. This operation increases $w(E[U^{{}^{\prime\prime}}])$}.
For scenarios where $\lvert C\rvert>k$, two operations exist. The operation that removes nodes from $C^{{}^{\prime\prime}}\setminus U^{\prime}$ further adds nodes to $U^{{}^{\prime\prime}}$, thereby increasing $w(E[U^{{}^{\prime\prime}}])$}. It remains to consider the operation, which greedily removes nodes from $C^{{}^{\prime\prime}}\cap U^{\prime}$ in $C^{{}^{\prime\prime}}$ and which decreases $w(E[U^{{}^{\prime\prime}}])$}.
We now define $U^{{}^{\prime\prime}}_{0}$ as the initial $U^{{}^{\prime\prime}}$ of this greedy procedure. As the previous two operations contribute additional nodes to $U^{{}^{\prime\prime}}$, we conclude that ${w(E[U^{{}^{\prime\prime}}_{0}])}\geq{w(E[U^{\prime}])}$. Similarly, we assign $C^{{}^{\prime\prime}}_{0}$ as the initial $C^{{}^{\prime\prime}}$ of the greedy procedure.
Let $U^{{}^{\prime\prime}}_{t}$ represent $U^{{}^{\prime\prime}}$ and $C^{{}^{\prime\prime}}_{t}$ represent $C^{{}^{\prime\prime}}$ at the $t$-th step of the greedy procedure. We observe that at the $t$-th step, the decrease in the sum of edge weights is at most $\frac{2{w(E[U^{{}^{\prime\prime}}_{t}])}}{\lvert C^{{}^{\prime\prime}}_{t}\rvert}$, as suggested by the pigeonhole principle. In the worst-case scenario, the sum of edge weights decreases $\lvert C^{{}^{\prime\prime}}_{0}\rvert-k$ times, until $\lvert C^{{}^{\prime\prime}}\rvert=k$. The value of $w(E[U^{{}^{\prime\prime}}])$} at this point is lower bounded by:
$\begin{split}&{w(E[U^{{}^{\prime\prime}}_{0}])}\left(1-\frac{2}{\lvert C^{{}^{\prime\prime}}_{0}\rvert}\right)\left(1-\frac{2}{\lvert C^{{}^{\prime\prime}}_{1}\rvert}\right)\ldots\left(1-\frac{2}{k+1}\right)\\ &={w(E[U^{{}^{\prime\prime}}_{0}])}\frac{(\lvert C^{{}^{\prime\prime}}_{0}\rvert-2)}{\lvert C^{{}^{\prime\prime}}_{0}\rvert}\frac{(\lvert C^{{}^{\prime\prime}}_{1}\rvert-2)}{\lvert C^{{}^{\prime\prime}}_{1}\rvert}\ldots\frac{k-1}{k+1}\\ &={w(E[U^{{}^{\prime\prime}}_{0}])}\frac{k(k-1)}{\lvert C^{{}^{\prime\prime}}_{0}\rvert(\lvert C^{{}^{\prime\prime}}_{0}\rvert-1)}.\end{split}$
Given that $\lvert C^{{}^{\prime\prime}}_{0}\rvert\leq\lvert C\rvert$ and ${w(E[U^{{}^{\prime\prime}}_{0}])}\geq{w(E[U^{\prime}])}$, the above expression is at least ${w(E[U^{\prime}])}\frac{k(k-1)}{\lvert C\rvert(\lvert C\rvert-1)}\geq{w(E[U^{\prime}])}\frac{k(k-1)}{\lvert C\rvert^{2}}$. Notice that we assume $k\in\Omega(n)$, and thus $1-\frac{1}{k}=1-\frac{1}{\Theta(n)}$. This concludes our proof. ∎
Recall that our goal is to compute the approximation ratio, which can be formulated as
$\max_{\gamma\in[0.1,5],\eta\in\left[\frac{1-\alpha+0.01}{\alpha},+\infty\right)}\min_{\mu\in[0,1]}\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma,\eta).$ (21)
Let us define $f_{\tau}(\mu,\gamma,\eta)=\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma,\eta)$. Notice that for $k$-Densify, according to the above lemma, when $\mu\leq\tau$, $\kappa_{\tau}(\mu)=1$. Hence, for the worst-case analysis, we only need to discuss the scenario when $\mu\in(\tau,1]$. Below, we give more details on computing the approximation ratio. We ignore the factor $(1-\frac{1}{\Theta(n)})$ in our computation as it only introduces a small multiplicative error in our result.
C.5.2 Computing the approximation ratio
We first find the stationary point of $f_{\tau}(\mu,\gamma,\eta)$ w.r.t. $\mu$ by solving
$\frac{\partial f_{\tau}(\mu,\gamma,\eta)}{\partial\mu}=0.$ (22)
Let this stationary point be $\mu_{0}$. We notice that $\mu_{0}$ can be written as a function of $\gamma$, $\eta$ and $\tau$. Since $\mu_{0}$ should be in the domain $\tau,1]$, we introduce the following step function $g_{\tau}(\gamma,\eta)$ to tackle the situation when $\mu_{0}$ is invalid (chosen outside of the domain):
[…]
For any given $\gamma$ and $\eta$, the minimization of the function $f_{\tau}(\mu,\gamma,\eta)$ w.r.t. $\mu$ can be written as the minimum of three constant functions. The two boundary cases when $\mu=\tau$ or $\mu=1$ are given by $f_{\tau}(\mu_{0},\gamma,\eta)$ and $f_{\tau}(1,\gamma,\eta)$, whereas $f_{\tau}(\mu_{0},\gamma,\eta)$ represents the case when $\mu\in[\tau,1]$. Thus, $\min_{\mu\in[\tau,1]}f_{\tau}(\mu,\gamma,\eta)=\min\left\{f_{\tau}(\tau,\gamma,\eta),f_{\tau}(1,\gamma,\eta),g_{\tau}(\gamma,\eta)\right\}$. Now we obtain:
$f(\tau)$ $=\max_{\gamma,\eta}\min_{\mu\in[\tau,1]}f_{\tau}(\mu,\gamma,\eta)$
$=\max_{\gamma,\eta}\min\left\{f_{\tau}(\tau,\gamma,\eta),f_{\tau}(1,\gamma,\eta),g_{\tau}(\gamma,\eta)\right\}.$
Then for any $\tau\in(0,\frac{1}{2}]$, we numerically select $\gamma$ and $\eta$ to maximize $f_{\tau}(\cdot)$ and $g_{\tau}(\cdot)$, and we obtain Figure 8.
Figure 8: Our approximation ratios for $k\text{-}{{\small D}\textsc{ensify}}$ where $\tau=\frac{k}{n}$. We compare the approximation ratio $f(\tau)$ with $\tau$ and $\tau^{2}$. Here, $f(\tau)$ is the worst case approximation ratio of our SDP-based algorithm for the $k$-Densify problem. The approximation ratio for DS-kkR is $f(\tau)\frac{1-c}{1+c}$.
In particular, if we set $\gamma=0.920$ and $\eta=1.65$, we get $f(\tau=0.5)>0.58$. Note if we set the same parameter as Feige and Langberg [2001] ($\gamma=3.87$ and $\eta=\frac{0.65}{3.87}$), we obtain the same approximation ratio as Feige and Langberg [2001] obtained for DkkS with $f(\tau=0.5)=0.517$.
C.6 Pseudocode for SDP-based algorithm for Max-Cut-kkR (in Section 5.1)
The pseudocode for the SDP-based algorithm for Max-Cut-kkR, as discussed in Section 5.1, is illustrated in Algorithm 4.
Algorithm 4 An SDP-based algorithm for Max-Cut-kkR
1:$(G(V,E,w),U,k)$, number of iterations NN
2:Step 1: Solve SDP. Solve the SDP (3) to obtain the solution $\bm{v}_{0},\ldots,\bm{v}_{n}$
3:for $i=1$ to NN do
4:  Step 2: Hyperplane Rounding. Obtain the indicator vector $\overline{\bm{x}}$
5:  Sample a unit vector $\bm{r}$ uniformly randomly from $\mathcal{S}_{n}$
6:  for all $i\in\{0,\ldots,n\}$ do
7:   if $(\bm{r}\cdot\bm{v}_{0})(\bm{r}\cdot\bm{v}_{i})\geq 0$ then
8:     $\overline{x}_{i}\leftarrow 1$
9:   else
10:     $\overline{x}_{i}\leftarrow-1$
11:   end if
12:  end for
13:  Set $C\leftarrow\{i\colon\overline{x}_{i}\neq x^{0}_{i}\}$
14:  Step three, Fix $C$: Set $C^{{}^{\prime\prime}}\leftarrow C$
15:  while $\lvert C^{{}^{\prime\prime}}\rvert\neq k$ do
16:   if $\lvert C^{{}^{\prime\prime}}\rvert<k$ then
17:     $u^{*}\leftarrow\arg\max_{u\in V\setminus C^{{}^{\prime\prime}}}\mathrm{cut}\!\left(U\triangle(C^{{}^{\prime\prime}}\cup\{u\})\right)$
18:     $C^{{}^{\prime\prime}}\leftarrow C^{{}^{\prime\prime}}\cup\{u^{*}\}$
19:   else if $\lvert C^{{}^{\prime\prime}}\rvert>k$ then
20:     $u^{*}\leftarrow\arg\max_{u\in C^{{}^{\prime\prime}}}\mathrm{cut}\!\left(U\triangle(C^{{}^{\prime\prime}}\setminus\{u\})\right)$
21:     $C^{{}^{\prime\prime}}\leftarrow C^{{}^{\prime\prime}}\setminus\{u^{*}\}$
22:   end if
23:  end while
24:end for
25:return $C^{{}^{\prime\prime}}$
C.7 Proof of Theorem 5.2
The pseudocode for our algorithm can be found in Algorithm 4. In the present section, we demonstrate that Algorithm 4 provides a constant factor approximation to the Max-Cut-kkR problem. Note that the first two steps (i.e., solving the SDP and hyperplane rounding) are identical to those of Algorithm 3. The third step (i.e., fixing $C$) is identical to the third step of Algorithm 2.
C.7.1 Fixing $C$
Again we let $\lvert C\rvert=\mu n$ and $k=\tau n$. We notice the greedy procedure (i.e., fixing $C$) that appears in Algorithm 4 is the same as the counterpart that appears in Algorithm 2. Thus, we can directly apply Lemma B.2 to derive a lower bound on the cut for fixing $C$, as in Lemma C.9.
Lemma C.9.
For Max-Cut-kkR, $\mathbf{SDP}\geq\min\{\frac{\tau^{2}}{\mu^{2}}(1-\frac{1}{\Theta(n)}),\frac{(1-\tau)^{2}}{(1-\mu)^{2}}(1-\frac{1}{\Theta(n)})\}\overline{\mathbf{SDP}}$. In particular, when $\mu>\tau$, the cut is lower bounded by $\frac{\tau^{2}}{\mu^{2}}(1-\frac{1}{\Theta(n)})\mathrm{cut}\!\left(U^{\prime}\right)$; when $\mu<\tau$, the cut is lower bounded by $\frac{(1-\tau)^{2}}{(1-\mu)^{2}}(1-\frac{1}{\Theta(n)})\mathrm{cut}\!\left(U^{\prime}\right)$.
Proof.
The proof is identical to Lemma B.2. ∎
C.7.2 Approximation ratio
Let $k=\tau n$ and let $\lvert C\rvert=\mu n$. Again, we analyze the approximation ratio of Max-Cut-kkR when $k\leq n/2$. By symmetry, we can obtain the approximation ratio of Max-Cut-kkR when $k>n/2$.
We start with the following corollary of Lemma C.4. Notice that the corollary slightly differs from Lemma C.4 in terms of introducing different random variables $Z_{1}$ and $Z_{2}$ for the two cases $\lvert C\rvert\geq k$ and $\lvert C\rvert<k$, respectively. This design helps to obtain better approximation ratios for Max-Cut-kkRby selecting better parameters for the two cases separately.
Corollary C.4.
Consider the constants $\gamma_{1},\gamma_{2}\in[0.1,5]$ and $\eta_{1},\eta_{2}\in\left[\frac{1-\alpha+0.01}{\alpha},+\infty\right)$. Define the random variables $Z_{1}$ and $Z_{2}$ as follows:
$\begin{cases}Z_{1}=\frac{\mathrm{cut}\!\left(U^{\prime}\right)}{W^{*}_{SDP}}+\gamma_{1}\eta_{1}\frac{n-\lvert C\rvert}{n-k}+\gamma_{1}\frac{\lvert C\rvert(2k-\lvert C\rvert)}{n^{2}},&\text{if }\lvert C\rvert>k,\\ Z_{2}=\frac{\mathrm{cut}\!\left(U^{\prime}\right)}{W^{*}_{SDP}}+\gamma_{2}\eta_{2}\frac{\lvert C\rvert}{k}+\gamma_{2}\frac{\lvert C\rvert(2k-\lvert C\rvert)}{n^{2}},&\text{if }\lvert C\rvert<k.\end{cases}$
If the hyperplane rounding is repeated at least $N=2\cdot\frac{1-p+\epsilon p}{\epsilon p}\log(\frac{1}{\epsilon})$ times, where $p$ is a constant with a lower bound of $0.0005$, then one of the following holds: either $\lvert C\rvert\geq k$ occurs at least $\frac{N}{2}$ times, and with probability at least $1-\epsilon$, $Z_{1}\geq(1-\epsilon)[\alpha+\gamma_{1}\eta_{1}\alpha+\gamma_{1}(\alpha(1-\tau)^{2}-1+2\tau)]$; or $\lvert C\rvert<k$ occurs at least $\frac{N}{2}$ times, and with probability at least $1-\epsilon$, $Z_{2}\geq(1-\epsilon)[\alpha+\gamma_{2}\eta_{2}\alpha+\gamma_{2}(\alpha(1-\tau)^{2}-1+2\tau)]$.
Next, we use Corollary C.4 to bound our approximation ratio by solving
$\max_{\gamma\in[0.1,5],\eta\in\left[\frac{1-\alpha+0.01}{\alpha},+\infty\right)}\min_{\mu\in[0,1]}\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma,\eta).$ (24)
Notice that according to Lemma C.9, $\kappa_{\tau}(\mu)=\frac{\tau^{2}}{\mu^{2}}(1-\frac{1}{\Theta(n)})$ when $\lvert C\rvert\geq k$ i.e., $\mu\geq\tau$, and $\kappa_{\tau}(\mu)=\frac{(1-\tau)^{2}}{(1-\mu)^{2}}(1-\frac{1}{\Theta(n)})$ when $\lvert C\rvert<k$ i.e., $\mu<\tau$. According to Corollary C.4,
$\lambda_{\tau}(\mu,\gamma_{1},\eta_{1})\geq\alpha+\gamma_{1}\eta_{1}\alpha+\gamma_{1}(\alpha(1-\tau)^{2}-1+2\tau)-\gamma_{1}\eta_{1}(1-\mu)\frac{1}{1-\tau}-\gamma_{1}\mu(2\tau-\mu)-\frac{1}{\Theta(n)},$ (25)
when $\mu\geq\tau$; and
$\lambda_{\tau}(\mu,\gamma_{2},\eta_{2})\geq\alpha+\gamma_{2}\eta_{2}\alpha+\gamma_{2}(\alpha(1-\tau)^{2}-1+2\tau)-\gamma_{2}\eta_{2}\frac{\mu}{\tau}-\gamma_{2}\mu(2\tau-\mu)-\frac{1}{\Theta(n)},$ (26)
when $\mu<\tau$.
We ignore the factor $(1-\frac{1}{\Theta(n)})$ in our computation as it only introduces a small multiplicative error in our result.
Case 1: $\lvert C\rvert\geq k$. This implies that $1\geq\mu\geq\tau$. In this domain of $\mu$, let us define the function $f^{1}_{\tau}(\mu,\gamma_{1},\eta_{1})=\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma_{1},\eta_{1})$. The approximation ratio we get from this case is
$f^{1}(\tau)=\max_{\gamma_{1}\in[0.1,5],\eta_{1}\in\left[\frac{1-\alpha+0.01}{\alpha},+\infty\right)}\min_{\mu\in[\tau,1]}f^{1}_{\tau}(\mu,\gamma_{1},\eta_{1}).$
Case 2: $\lvert C\rvert<k$. This implies $\tau>\mu>0$. In this domain of $\mu$, let us define the function $f^{2}_{\tau}(\mu,\gamma_{2},\eta_{2})=\kappa_{\tau}(\mu)\lambda_{\tau}(\mu,\gamma_{2},\eta_{2})$. The approximation ratio we get from this case is
$f^{2}(\tau)=\max_{\gamma_{2}\in[0.1,5],\eta_{2}\in\left[\frac{1-\alpha+0.01}{\alpha},+\infty\right)}\min_{\mu\in(0,\tau]}f^{2}_{\tau}(\mu,\gamma_{2},\eta_{2}).$
Combining Case 1 and Case 2, consider $1\geq\mu>0$. We define the function
$h(\tau)=\min\left\{f^{1}(\tau),f^{2}(\tau)\right\},$
and note that $h(\tau)$ is the approximation ratio we get in the end. We plot $h(\tau)$ in Figure 9.
Figure 9: Our approximation ratios for Max-Cut-kkR, where $\tau=\frac{k}{n}$. We compare the approximation ratio $h(\tau)$ of our SDP-based algorithm with $\min\{\tau,1-\tau\}$ and $\min\{\tau^{2},(1-\tau)^{2}\}$.
Appendix D Omitted content and proofs of Section 6
In this section, we elaborate on the sum-of-squares (SoS) algorithm from Section 6. In particular, we apply the method proposed by  Raghavendra and Tan [2012]. In detail, we will explain how to relax Max-Cut-kkR into the Lasserre SDP hierarchy (or, Lasserre hierarchy), how to obtain globally uncorrelated solutions to the Lasserre hierarchy, and how to apply a rounding scheme to the solutions so that the cardinality constraint is almost satisfied.
For a detailed description and properties of the Lasserre SDP hierarchy, we refer the readers to [Lasserre, 2002; Rothvoß, 2013]. For more technical details in this section, we refer the readers to [Austrin et al., 2016; Raghavendra and Tan, 2012; Chlamtac and Tulsiani, 2012].
The remainder of this section is structured as follows. Section D.1 briefly introduces properties of the Lasserre hierarchy and mutual information to be applied in the following analysis. Section D.2 formulates a degree $d$ Lasserre hierarchy program for Max-Cut-kkR. Section D.3 describes how to obtain a decorrelated solution via conditioning. Section D.4 formulates the degree $2$ Lasserre hierarchy and the hyperplane rounding algorithm we apply. Section D.5 proves Lemma 6.2 that we presented in the main content.
D.1 Preliminaries
D.1.1 Notation
Let $G=(V,E,W)$ be an undirected weighted graph, and let $\bm{x}^{0}\in\left\{-1,1\right\}^{V}$ encode the initial partition, i.e., $(U,\overline{U})$. The $i$-th entry of $\bm{x}^{0}$ is denoted by $x^{0}_{i}$. Let $S\subseteq V$, and $\mu_{S}$ be a (pseudo)-distribution on $\{-1,1\}^{S}$. We will elaborate on its properties in Section D.1.2. Let $\bm{x}\in\left\{-1,1\right\}^{V}$ be a random vector that encodes a random partition. We use $\bm{x}_{S}$ to indicate the partition constrained on $S$. Moreover, for $\xi\in\{-1,1\}^{S}$, we denote by $\Pr_{\mu_{S}}(\bm{x}_{S}=\xi)$ the probability of assigning $\bm{x}_{S}$ with $\xi$ under the distribution $\mu_{S}$.
Let $k\leq\lvert V\rvert$ be the number of refinements and $k=\tau n$. We let $\varsigma=(n-2k)/n,\varsigma\in(-1,1)$. Let us fix $\varepsilon>0$ to be a small constant, and let $\delta=(\varepsilon/3)^{96}$. Let $d^{\prime}=\lceil 1+\frac{1}{\delta}\rceil$. We will construct a degree $d=d^{\prime}+2$ Lasserre hierarchy.
D.1.2 Properties of the Lasserre hierarchy
Consistent local distribution. Any Lasserre hierarchy configuration can be written in terms of local distributions [Lasserre, 2002]. In particular, in the degree $d$ Lasserre hierarchy, for each set $S\subseteq V,|S|\leq d$, there is a distribution $\mu_{S}$ on $\left\{-1,1\right\}^{S}$. Furthermore, Lasserre hierarchy ensures that for any two such probability distributions, one on $\{-1,1\}^{S}$, denoted as $\mu_{S}$, and one on $\{-1,1\}^{T}$, denoted as $\mu^{\prime}_{T}$, their marginal distribution on $\{-1,1\}^{S\cap T}$ is the same: $\mu_{S\cap T}=\mu^{\prime}_{S\cap T}$. In other words, local probability distributions are consistent.
Semi-definiteness property. Local distributions of the Lasserre hierarchy also satisfy positive semi-definiteness properties. In the degree $d$ Lasserre hierarchy, for each $S,T\subseteq V,|S\cup T|\leq d,$ and $\xi_{1}\in\left\{-1,1\right\}^{S},\xi_{2}\in\left\{-1,1\right\}^{T}$, there exist vectors $\bm{v}_{S,\xi_{1}},\bm{v}_{T,\xi_{2}}$ such that
$\Pr_{\mu_{S\cup T}}(\bm{x}_{S}=\xi_{1},\bm{x}_{T}=\xi_{2})=\bm{v}_{S,\xi_{1}}\cdot\bm{v}_{T,\xi_{2}},$
where $\bm{x}_{T},\bm{x}_{S}$ are random variables taking values in $\left\{-1,1\right\}^{S},\left\{-1,1\right\}^{T}$, respectively, and $\cdot$ denotes the dot product of two vectors.
Conditioning of assignments. One property that we will use in our algorithm is that the local probability distributions can be conditioned. In particular, any event $\bm{x}_{S}=\zeta$ can be conditioned on an an event $\bm{x}_{U}=\xi$ given that $\lvert S\cup U\rvert\leq d$ and $\Pr_{\mu_{S\cup U}}(\bm{x}_{U}=\xi)\neq 0$.
After the conditioning, we obtain a new Lasserre hierarchy of degree d|U|d-|U| with local distributions $\left\{\bar{\mu}_{S}\right\}_{S\subseteq V,|S|\leq d-|U|}$. In particular,
$\begin{split}\Pr_{\bar{\mu}_{S}}\left(\bm{x}_{S}=\zeta\right)=\Pr_{\mu_{S\cup U}}\left(\bm{x}_{S}=\zeta\mid\bm{x}_{U}=\xi\right)=\frac{\bm{v}_{S,\zeta}\cdot\bm{v}_{U,\xi}}{\bm{v}_{U,\xi}\cdot\bm{v}_{U,\xi}}.\end{split}$
D.1.3 Mutual information
Next, we illustrate how to use conditioning to reduce the correlation between fractional solutions of the vertices, allowing us to round values of different vertices almost independently on average. We need to use mutual information to measure the correlation between vertices.
Definition D.1.
Let $x_{i},x_{j}$ be two random variables taking values in $\left\{-1,1\right\}$, which indicates the partition of node $i$ and $j$. Let $S\supseteq\left\{i,j\right\},U\ni i,T\ni j$ in place of $\left\{i,j\right\},\left\{i\right\},\left\{j\right\}$. The mutual information between $x_{i}$ and $x_{j}$ is then defined as
$I(x_{i},x_{j})=\sum_{a,b\in\{-1,1\}}\Pr_{\mu_{S}}(x_{i}=a,x_{j}=b)\cdot\log\left(\frac{\Pr_{\mu_{S}}(x_{i}=a,x_{j}=b)}{\Pr_{\mu_{U}}(x_{i}=a)\Pr_{\mu_{T}}(x_{j}=b)}\right).$
This follows from the fact that probability distributions are consistent under intersections, and hence we can take $S\cap\left\{i,j\right\}$ to obtain the distribution on $\left\{i,j\right\}$, etc.
We will aim to get
$\operatorname*{\mathbb{E}}_{i,j\sim V}\left[I(x_{i},x_{j})\right]\leq\delta.$
This goal will be achieved by conditioning and utilizing the observation from Lemma D.1. In particular, Lemma D.1 implies that the following equation holds:
$\min_{\begin{subarray}{c}\{i_{1},\ldots,i_{t}\}\subseteq V,t\leq d^{\prime}\\ \{a_{i_{1}},\ldots,a_{i_{t}}\}\in\{-1,1\}^{\{i_{1},\ldots,i_{t}\}}\end{subarray}}\operatorname*{\mathbb{E}}_{i,j\sim V}\left[I(x_{i},x_{j}\mid x_{i_{1}}=a_{i_{1}},\ldots,x_{i_{t}}=a_{i_{t}})\right]\leq 1/(d^{\prime}-1).$ (27)
Lemma D.1 (Lemma 4.5 in Raghavendra and Tan [2012]).
There exists $t\leq d^{\prime}$ such that
$\operatorname*{\mathbb{E}}_{i_{1},\ldots,i_{t}\sim V}\operatorname*{\mathbb{E}}_{i,j\sim V}\left[I(x_{i},x_{j}\mid x_{i_{1}},\ldots,x_{i_{t}})\right]\leq 1/(d^{\prime}-1).$
This lemma holds true for any Lasserre hierarchy over a binary [^5] The original lemma was shown for nonbinary alphabets (alphabets of size $q$), with the right-hand side containing an additional $\log_{2}q$ factor. alphabet.
D.2 Lasserre hierarchy program for Max-Cut-kkR
The Lasserre hierarchy relaxation of degree $d$ for Max-Cut-kkR is as follows:
$\max$ $\sum_{i<j}W_{ij}\Pr_{\mu_{\{i,j\}}}(x_{i}\neq x_{j})$ (28)
s.t. $\sum_{j=1}^{n}\Pr_{\mu_{S\cup\{j\}}}(x_{j}=x_{j}^{0}\cap\bm{x}_{S}=\xi)=(n-k)\Pr_{\mu_{S}}(\bm{x}_{S}=\xi),$
$\forall S\subseteq V,\xi\in\{-1,1\}^{S},|S\cup\{j\}|<d.$
We can show that Equation (28) is indeed a relaxation of Max-Cut-kkR. Let $\hat{\bm{x}}$ be any feasible solution of Max-Cut-kkR, and notice that it satisfies the constraint $\sum_{i}x_{i}^{0}\hat{x}_{i}=n-2k$. Now we independently assign that $\Pr_{\mu_{S\cup\{j\}}}(x_{j}=\hat{x}_{j})=1$, and check whether this assignment satisfies the constraint. Notice that for any $\bm{x}_{S}$ and $\xi$, if $\Pr_{\mu_{S\cup\{j\}}}(\bm{x}_{S}=\xi)=0$, the constraint trivially holds. Otherwise, we condition the probability on $\bm{x}_{S}=\xi$, and we get
$\sum_{j=1}^{n}\Pr_{\mu_{S\cup\{j\}}}(x_{j}=x_{j}^{0}\mid\bm{x}_{S}=\xi)$ $=\sum_{j=1}^{n}\Pr_{\mu_{S\cup\{j\}}}(\hat{x}_{j}=x_{j}^{0}\mid\bm{x}_{S}=\xi)$
$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{n}\Pr_{\mu_{S\cup\{j\}}}(\hat{x}_{j}=x_{j}^{0})$
$\stackrel{{\scriptstyle(b)}}{{=}}n\frac{(n-k)}{n}=n-k.$
In the above formulation, $(a)$ holds as the assignment is independent, and $(b)$ holds by the constraint from Equation 2 that $\sum_{i}x_{i}^{0}\hat{x}_{i}=n-2k$.
Notice that an equivalent SDP can be formulated as follows. For every $S\subseteq V,|S\cup\{j\}|\leq d$, and $\xi\in\{-1,1\}^{S}$, we introduce a variable $\bm{v}_{S,\xi}$. Then we obtain:
$\max$ $\sum_{i<j}W_{ij}(\bm{v}_{i,1}\cdot\bm{v}_{j,-1}+\bm{v}_{i,-1}\cdot\bm{v}_{j,1})$ (29)
s.t. $\sum_{j=1}^{n}x_{j}^{0}\bm{v}_{j,1}\cdot\bm{v}_{S,\xi}-\sum_{j=1}^{n}x_{j}^{0}\bm{v}_{j,-1}\cdot\bm{v}_{S,\xi}=(n-2k)\bm{v}_{S,\xi}\cdot\bm{v}_{S,\xi},$
$\forall S\subseteq V,\xi\in\{-1,1\}^{S},|S\cup\{j\}|<d.$
Since this is a semidefinite relaxation with a polynomial ($n^{O(d)}$) number of constraints, we can obtain the optimal value of this program in polynomial time up to precision $\delta>0$. Let us write $\mathbf{Sol}$ to denote the value of this solution and $\operatorname{\bf Opt}$ to denote the optimal value Max-Cut-kkR. We also use $\left\{\bm{v}_{S}\right\}_{S\subseteq V,|S|\leq d}$ to denote the vectors $\bm{v}_{S}$ under which the value $\mathbf{Sol}$ is attained, and hence we have
$(1-\delta)\operatorname{\bf Opt}\leq\sum_{i<j}W_{ij}(\bm{v}_{i,1}\cdot\bm{v}_{j,-1}+\bm{v}_{i,-1}\cdot\bm{v}_{j,1})=\mathbf{Sol}.$
D.3 Obtaining a “decorrelated” solution via conditioning
After we solve Equation (28), we obtain the local distributions of $\bm{x}$, which are consistent on the marginal distributions up to $d$ entries. We will apply a conditioning process in order to obtain a Lasserre hierarchy such that the mutual information of each pair of entries is bounded, i.e., $\operatorname*{\mathbb{E}}_{i,j\sim V}\left[I(x_{i},x_{j})\right]\leq\delta$. In particular, a step in the conditioning the procedure works as follows:
  • Pick a vertex $i$ uniformly at random, and fix its value to either $1$ or 1-1, according to its marginal distribution $\mu_{\{i\}}$. Update the probability distributions $\{\mu_{S}\}_{S\subseteq V}$ according to this fixing, as discussed in Subsection D.1.2.
Let us observe that the objective value of Equation (28) does not change in expectation. Furthermore, if we denote $\tilde{\bm{v}}_{j,-1},\tilde{\bm{v}}_{j,1},\tilde{\bm{v}}_{S,\xi}$ the vectors obtained after the conditioning, then in expectation we obtain that
$\sum_{j=1}^{n}x_{j}^{0}\tilde{\bm{v}}_{j,1}\cdot\tilde{\bm{v}}_{S,\xi}-\sum_{j=1}^{n}x_{j}^{0}\tilde{\bm{v}}_{j,-1}\cdot\tilde{\bm{v}}_{S,\xi}=(n-2k)\tilde{\bm{v}}_{S,\xi}\cdot\tilde{\bm{v}}_{S,\xi}$
is satisfied.
Next, we apply Lemma D.1 to obtain Lemma D.2, which is analogous to Lemma 4.6 from Raghavendra and Tan [2012]. The difference is that our problem Max-Cut-kkR has a different constraint; hence, a different analysis on the lower bound on the objective value is needed in order to apply the Markov inequality.
Lemma D.2.
There exists an algorithm running in time
$O\left(n^{2}d^{\prime}\left(\frac{en}{d^{\prime}}\right)^{d^{\prime}}2^{d^{\prime}}\right),$
which, given a Lasserre solution outlined in Equation (28) of degree $d=d^{\prime}+2$ with value $\mathbf{Sol}$, finds a degree $2$ Lasserre solution with objective value $\mathbf{Sol}(1-\delta)$, and which satisfies $\operatorname*{\mathbb{E}}[I(x_{i},x_{j})]\leq\delta$.
Proof.
Using the probabilistic method, we first show that with some probability, conditioning of variables does not reduce the objective value by more than $\delta\mathbf{Sol}$. Then, we show that with high probability, conditioning satisfies the inequality
$\operatorname*{\mathbb{E}}[I(x_{i},x_{j})]\leq\delta.$
The randomness in the expression above comes from the random choices in our conditioning algorithm. Then, by union bound, we can conclude that there is a conditioning with value at least $\mathbf{Sol}-\delta\mathbf{Sol}$ and with low average mutual information. We then conclude the lemma by discussing how this conditioning can be found efficiently in the time described in the lemma statement.
Let us now provide details. Let us denote with $W\in\mathbb{R}_{+}$ the value $W=\sum_{i<j}W_{ij}$. First, consider the conditioning procedure outlined in this section. Since we are sampling the variables according to their marginal distributions, the expected value of the program (28) remains $\mathbf{Sol}$. Let us use $X$ to denote a random variable whose value is the value of the Lasserre hierarchy after $t$ conditioning steps, where $t$ is the same number as the number from Lemma D.1. We can then calculate
$\Pr\left(X<(1-\delta)\operatorname*{\mathbb{E}}[X]\right)=\Pr\left(W-X\geq W-(1-\delta)\operatorname*{\mathbb{E}}[X]\right).$
Now, since $X\in[0,W]$, by Markov’s inequality we have
$\Pr\left(X<(1-\delta)\operatorname*{\mathbb{E}}[X]\right)\leq\frac{W-\operatorname*{\mathbb{E}}[X]}{W-(1-\delta)\operatorname*{\mathbb{E}}[X]}=\frac{1-\operatorname*{\mathbb{E}}[X]/W}{1-(1-\delta)\operatorname*{\mathbb{E}}[X]/W}.$ (30)
Let us use $\zeta=\operatorname*{\mathbb{E}}[X]/W$. In order to find a suitable upper bound on the expression above, we first prove a lower bound on $\zeta$.
For this, we consider a feasible solution to Equation (28): for each $i\in V$, we independently assign $x_{i}^{0}$ to vertex $i$ according to probability $\Pr_{\mu_{i}}(x_{i}=x_{i}^{0})=\frac{n-k}{n}=(1+\varsigma)/2$, and $x_{i}^{0}$ otherwise. Let us now give a lower bound for $\operatorname*{\mathbb{E}}[X]$, by giving a lower bound of a term
$\Pr_{\mu_{\{i,j\}}}(x_{i}\neq x_{j})=\frac{1-\varsigma^{2}}{2},$
according to the assignment. Going back to Equation (30), this shows that $\zeta=\operatorname*{\mathbb{E}}[X]/W$ satisfies $\zeta\geq(1-\varsigma^{2})/2$. Hence, we can write
$\Pr\left(X<(1-\delta)\operatorname*{\mathbb{E}}[X]\right)\leq\frac{1-\zeta}{1-(1-\delta)\zeta}\leq\frac{1}{1+\delta\frac{1-\varsigma^{2}}{1+\varsigma^{2}}}.$
Let us denote with $I$ the random variable $I=I(X_{i},X_{j}\mid X_{i_{1}},\ldots,X_{i_{t}})$. By Markov’s inequality
$\Pr\left(I\geq\sqrt{\frac{1}{d^{\prime}-1}}\right)\leq\frac{\operatorname*{\mathbb{E}}[I]}{\sqrt{\frac{1}{d^{\prime}-1}}}\leq\sqrt{\frac{1}{d^{\prime}-1}},$
and therefore the probability of the randomized algorithm not finding a fixing for which $\operatorname*{\mathbb{E}}[I_{i,j}]\leq\sqrt{\frac{1}{d^{\prime}-1}}$ is at most $\sqrt{\frac{1}{d^{\prime}-1}}$. Hence, the probability of the algorithm failing is at most
$\begin{split}\frac{1}{1+\delta\cdot(1-\varsigma^{2})/(1+\varsigma^{2})}+\sqrt{\frac{1}{d^{\prime}-1}}=\frac{1}{1+\delta\cdot(1-\varsigma^{2})/(1+\varsigma^{2})}+\frac{1}{4}\cdot\delta(1-\varsigma^{2})<1,\end{split}$
and therefore by union bound there is a fixing which satisfies $\operatorname*{\mathbb{E}}[I_{i,j}]\leq\delta$ and with value at least $\mathbf{Sol}(1-\delta)$.
Such a fixing, i.e., computing Equation (27), can be found using brute-force search. The size of the space of all possible fixings can be calculated as follows. First, we observe that there [^6] Since we do not know the value $t$ before hand, we need to test all $i=1,\ldots,d^{\prime}$. are $\sum_{i=0}^{d^{\prime}}{n\choose i}$ possible variables. Furthermore, once we fix $i$ variables, the space of possible values these variables can be assigned to is of size $2^{i}$. Hence, using the upper bound ${n\choose i}\leq(ne/i)^{i}$, the size of this space is at most
$d^{\prime}\left(\frac{en}{d^{\prime}}\right)^{d^{\prime}}2^{d^{\prime}}.$
Since calculating mutual information takes time $O(n^{2})$, this gives us the final running time of our algorithm. ∎
D.4 Description of the hyperplane rounding algorithm
In this section, we will describe the hyperplane rounding algorithm to the solution of the degree $2$ Lasserre hierarchy, which is obtained by using the conditioning outlined in the proof of Lemma D.2. Let us first write the formulation of the degree $2$ Lasserre hierarchy in Equation (31):
$\sum_{i<j}W_{ij}(\bm{v}_{i,1}\cdot\bm{v}_{j,-1}+\bm{v}_{i,-1}\cdot\bm{v}_{j,1})\geq\mathbf{Sol}(1-\delta),\quad\textrm{and}$ (31)
$\sum_{j=1}^{n}x_{j}^{0}\bm{v}_{j,1}\cdot\bm{v}_{i,1}-\sum_{j=1}^{n}x_{j}^{0}\bm{v}_{j,-1}\cdot\bm{v}_{i,1}=(n-2k)\bm{v}_{i,1}\cdot\bm{v}_{i,1},\quad\textrm{and}$
$\sum_{j=1}^{n}x_{j}^{0}\bm{v}_{j,1}\cdot\bm{v}_{i,-1}-\sum_{j=1}^{n}x_{j}^{0}\bm{v}_{j,-1}\cdot\bm{v}_{i,-1}=(n-2k)\bm{v}_{i,-1}\cdot\bm{v}_{i,-1}.$
If we define $\bm{v}_{i}\coloneqq\bm{v}_{i,-1}-\bm{v}_{i,1}$, and $\bm{v}_{0}\coloneqq\bm{v}_{i,-1}+\bm{v}_{i,1}$, the solution presented in Equation (31) implies a solution that satisfies the following constraints:
$\begin{split}\frac{1}{2}&\sum_{i<j}W_{ij}(1-\bm{v}_{i}\cdot\bm{v}_{j})\geq\mathbf{Sol}\cdot(1-\delta),\quad\textrm{and}\\ &\sum_{i\in V}x_{i}^{0}\bm{v}_{i}\cdot\bm{v}_{0}=\tau n.\end{split}$
Next, we describe the hyperplane rounding procedure, which differs from the ones in the other parts of our paper. The procedure is essentially the same as used by Raghavendra and Tan [2012, Section 5.4], with a mild difference in how we set the bias $\kappa_{i}$. Here we set $\kappa_{i}\coloneqq\bm{v}_{i}\cdot\bm{v}_{0}$. The goal is to ensure that the cardinality constraint in Max-Cut-kkR holds in expectation.
Let $\bar{\bm{v}}_{i}$ be defined as
[…]
Let us use $\varPhi$ to denote the cumulative distribution function of a single variable Gaussian distribution, and let us define the assignment $\{\bar{x}_{i}\}_{i=1}^{n}$ as
$\begin{split}g\sim\textrm{standard $n$-dimensional Gaussian vector},\\ \bar{x}_{i}=\begin{cases}-1,&\textrm{if }\bar{\bm{v}}_{i}\cdot g\leq\varPhi^{-1}\left(\frac{1-\kappa_{i}}{2}\right),\\ 1,&\textrm{otherwise.}\end{cases}\end{split}$
Notice that $\bar{\bm{x}}$ encodes the partition after the hyperplane rounding. We remark that that $\mathbb{E}\left[\bar{x_{i}}\right]=\kappa_{i}$. A computer-assisted proof shows that rounding the vectors $\bm{v}_{i}$ incurs a loss of at most $0.858$, as claimed in Raghavendra and Tan [2012, Section 5.4].
We present our main result of this hyperplane rounding procedure in Lemma D.3.
Lemma D.3.
After the hyperplane rounding procedure, it holds that:
  1. (a)
    $\frac{1}{2}\sum_{i<j}W_{ij}(1-\bar{x}_{i}\bar{x}_{j})\geq 0.858\frac{1}{2}\sum_{i<j}W_{ij}(1-\bm{v}_{i}\cdot\bm{v}_{j})$, and
  2. (b)
    $\mathbb{E}\left[\sum_{i\in V}x_{i}^{0}\bar{x}_{i}\right]=\tau n$.
Denote by $\overline{\mathbf{SoS}}$ the cut induced by the partition $\bar{\bm{x}}$ encodes. Then Part (a) of the lemma implies that $\overline{\mathbf{SoS}}\geq 0.858\mathbf{Sol}(1-\delta)$, and Part (b) implies that, in expectation, the cardinality constraint of Max-Cut-kkR is satisfied.
Lemma D.3 provides us with a cut that approximates the optimal solution well. However, the number of refinements satisfies the cardinality constraints only in expectation. In particular, $\sum_{i\in V}x_{i}^{0}\cdot\bar{x}_{i}=(n-2k)$ might do not hold.
In the last step we show that $\sum_{i\in V}x_{i}^{0}\cdot\bar{x}_{i}$ is very close to (n2k)(n-2k). Hence, we can change the signs of a very small number of $\bar{x}_{i}$ to obtain the final assignment $x_{i}$ for which the value of the objective function does not change much.
D.5 Obtaining a rounded solution which almost satisfies constraints
In the previous section, we proved that for our rounding procedure it holds that $\mathbb{E}\left[\sum_{i\in V}x_{i}^{0}\bar{x}_{i}\right]=\tau n$. In this section, we will show that $\sum_{i\in V}x_{i}^{0}\cdot\bar{x}_{i}$ is indeed very close to n2kn-2k, by bounding its variance.
Lemma D.4.
$\operatorname*{\mathbb{E}}_{i,j\in V}[I(x_{i},x_{j})]\leq\delta$ implies that
$\operatorname*{Var}\left[\sum_{i\in V}x_{i}^{0}\bar{x}_{i}\right]\leq C|V|^{2}\delta^{1/12},$
after we apply the hyperplane rounding procedure according to the above section, where $C$ is some constant.
Partial proof.
We only highlight the parts different from Raghavendra and Tan [2012, Theorem 5.6]. The differences arise as we consider the initial partition encoded as $\bm{x}_{0}$. The remaining part of the analysis is the same. We obtain:
$\begin{split}\operatorname*{Var}\left[\sum_{i\in V}x_{i}^{0}\bar{x}_{i}\right]&=\sum_{i,j\in V}(\mathbb{E}[\bar{x}_{i}x_{i}^{0}\bar{x}_{j}x_{j}^{0}]-\mathbb{E}[\bar{x}_{i}x_{i}^{0}]\mathbb{E}[\bar{x}_{j}x_{j}^{0}])\\ &\leq\sum_{i,j\in V}4\max_{a\in\{-1,1\},b\in\{-1,1\}}\Pr(\bar{x}_{i}=a,\bar{x}_{j}=b)-\Pr(\bar{x}_{i}=a)\Pr(\bar{x}_{i}=b)\\ &=\sum_{i,j\in V}\mathcal{O}\left(\sqrt{I(\bar{x}_{i},\bar{x}_{j})}\right)\\ &\leq C|V|^{2}\delta^{1/12}.\end{split}$
The last inequality follows directly from the analysis of Raghavendra and Tan [2012, Theorem 5.6]. ∎
Hence, by Chebyshev’s inequality, we have that
$\Pr\left(\left|\sum_{i\in V}x_{i}^{0}\bar{x}_{i}-(n-2\cdot k)\right|\geq\delta^{1/48}|V|\right)\leq\frac{|V|^{2}C(\delta^{1/12})}{\delta^{1/24}|V|^{2}}=C\delta^{1/24}.$
In other words, with probability $1-C\delta^{1/24}$ the values $\bar{x}_{i}$ satisfy
$\left|\sum_{i\in V}x_{i}^{0}\bar{x}_{i}-(n-2k)\right|\leq\delta^{1/48}|V|,$ (32)
Now we are ready to prove the main Lemma that we leave in the main content. See 6.2
Proof.
The coefficient $0.858(1-\delta)$ comes from the rounding algorithm described in the previous subsection and obtains a rounded solution of value at least $0.858$ from the SDP value, in expectation. We need to repeat rounding $\Omega(\log(1/\delta))$ times to ensure that with a high probability we obtain a rounded solution which incurs a loss of at most $0.858-\delta$. The other $(1-\delta)$-factor comes from the choice of precision when we solve the SDP formulation. Furthermore, for a single rounding step we have that
$\left|\sum_{i\in V}x_{i}^{0}\bar{x}_{i}-(n-2k)\right|\leq\delta^{1/48}|V|,$ (33)
with probability $1-C\delta^{1/24}$. Hence, for sufficiently small $\delta=(\varepsilon/3)^{96}$ by the union bound we can ensure that the properties stated in the lemma hold. ∎
Appendix E Additional experiments for DS-kkR
In this section, we present further details of our experiments. We run the experiments on a machine with an Intel Xeon Processor E5 2630 v4 and 64GB RAM. Our algorithms are implemented in Python. We use networkX for generating synthetic graphs, and for reading the real-world datasets. We use Mosek to solve the semidefinite programs.
E.1 Datasets
The first four datasets comprise Wikipedia politician page networks, and to obtain them we follow the method of Neumann and Peng [2022] with minor changes: instead of constructing edges according to the hyperlinks in the whole Wikipedia page, we only constructing edges using the hyperlinks that appear in the articles. By doing so, we avoid constructing edges simply because the two politicians belong to the same category. These networks represent politicians and party activities from different countries, annotated according to their respective parties. We select each graph based on the country and choose all the nodes belong to one party as the vertices of the initial subgraph.
The second type of datasets we utilize are obtained from SNAP networks [Leskovec and Krevl, 2014] with ground-truth communities. Specifically, we select SNAP-DBLP 777com-DBLP: https://snap.stanford.edu/data/com-DBLP.html, SNAP-Amazon 888com-Amazon: https://snap.stanford.edu/data/com-Amazon.html, and SNAP-Youtube 999com-Youtube: https://snap.stanford.edu/data/com-Youtube.html. For these datasets, we always choose the largest community as the initial subgraph.
We also use synthetic datasets generated from the Stochastic Block Model (SBM) with planted communities, namely SBM-Balanced, SBM-SparseSubg, and SBM-DenseSubg. Each graph has four communities. We generate SBM-SparseSubg and SBM-DenseSubg using the same parameters for generating the SBM, and they consist of three sparse communities and one dense community (see below for the concrete parameter values we picked). They only differ in their initial subgraphs. In SBM-SparseSubg, we use a sparse community as the initial subgraph, while in SBM-DenseSubg, we use the dense community as the initial subgraph. For SBM-Balanced, all four communities have equal density. With this configuration, the densest subgraph for SBM-Balanced is the whole graph, while for SBM-SparseSubg and SBM-DenseSubg, the denset subgraph is the planted dense community.
We elaborate on the parameters of the three SBM graphs. In each dataset, we establish four communities of equal size. For SBM-Balanced, all four communities are created with an edge probability of $0.3$, and the inter-community edge probabilities are set to $0.1$.
In contrast, for the SBM-SparseSubg and SBM-DenseSubg datasets, we create one dense community with edge probabilities $0.8$, alongside three sparse communities with edge probabilities $0.2$. The edge probability between communities for these two datasets remains at $0.1$.
Concerning the initial subgraph, we adopt a different community for each dataset. For SBM-Balanced, we choose any one of the planted communities. We select a sparse community for the SBM-SparseSubg dataset. Conversely, for SBM-DenseSubg, the dense community serves as our initial subgraph.
E.2 Additional experimental results for DS-kkR
E.2.1 Performance when initializing with ground-truth subgraphs with $k$ nodes removed
In Figure 10, we present additional experimental results when $k$ nodes were removed from a ground-truth subgraph.
SBM-SparseSubg SBM-DenseSubg SBM-Balanced
Figure 10: Relative increase of density for varying values of $k$. We initialize $U$ as a ground-truth subgraph and then remove $k$ nodes uniformly at random.
E.2.2 Performance when initializing with ground-truth subgraphs
In Table 3 and Figure 11, we present additional experimental results when $U$ is chosen a ground-truth subgraph.
SBM-SparseSubg SBM-DenseSubg SBM-Balanced
Figure 11: Relative increase of density for varying values of $k$. We initialized $U$ as a ground-truth subgraph.
E.2.3 Running time dependency on $k$
In Figure 12 we present the running time on synthetic datasets. Notice that for those synthetic datasets, we set a different range on $k$, i.e., $k\in[10,50]$. This is because it gives us a larger range on $k$ compared to $k\in[2\%n_{0},10\%n_{0}]$, since $n_{0}$ for those synthetic datasets is $250$. We notice that SDP, B: SDP and SQD are slower on SBM-SparseSubg than on SBM-DenseSubg. We do not get results for SQD on SBM-Balanced. In addition, SDP is always slower than B: SDP, since SDP contains more constraints.
SBM-Balanced SBM-DenseSubg SBM-SparseSubg
Figure 12: Running time for varying values of $k$. We initialize $U$ as a ground-truth subgraph and then remove $k$ nodes uniformly at random. We set $k=10,20,30,40,50$.
E.2.4 Running time dependency on $n$
We vary the number of nodes in the synthetic datasets SBM-DenseSubg and SBM-SparseSubg. Figure 13 shows the algorithms’ running times for different values of $n$. The running times of SDP and B: SDP increase quadratically with $n$ due to the high cost of solving the SDP, making these algorithms less scalable.
(a) SBM-DenseSubg (b) SBM-SparseSubg
Figure 13: Running time for varying values of $n$. We initialize $U$ as a ground-truth subgraph and then remove $k=10\%n_{0}$ nodes uniformly at random. We set n=1 000,1 250,1 500,1 750,2 000n=1\,000,1\,250,1\,500,1\,750,2\,000.
E.3 An analysis on SBM-SparseSubg
To better understand our experimental results on the SBM-SparseSubg dataset, we present an analysis of two naïve baselines. In particular, we consider SBM-SparseSubg where we initialize $U$ as a ground-truth subgraph and then remove $k$ nodes uniformly at random. In Figure 14 we plot the expected increase in density by two naïve baselines: We present the first baseline in blue, which directly recovers $k$ removed nodes (i.e., if $k=\lvert U\rvert$ it does exactly the same as Init). We present the second baseline in orange, which selects $k$ nodes from a dense community uniformly at random.
We observe that when $k\leq 84$, the baseline which picks nodes from the sparse community exhibits better performance. However, for larger values of $k$, the algorithm that select nodes from the dense community performs better. This figure can serve as a reference for us in understanding the selection of nodes by the actual algorithms.
Figure 14: On dataset SBM-SparseSubg. We first uniformly randomly remove $k$ nodes from the sparse community. We compute the expected relative increase on density by recovering $k$ nodes from the sparse community, or by select $k$ nodes from the dense community. When $k=84$ those two are equal.
In particular, a connection may be drawn between SBM-SparseSubg in Figure 10 and Figure 14. Upon analyzing the performances of SDP, B: SDP, and B: Peel, we find that when $k$ is small, their performances mirror that of the recovery of the sparse community, in the sense that their performances are closely aligned with Init. However, as $k$ increases, the performance of these three algorithms tends to resemble the selection of $k$ nodes from the dense community.
In contrast, Greedy always recovers the sparse community. Consequently, this algorithm does not exhibit a better performance when $k$ is large.
E.4 A discussion of SQD’s performance
Next, we provide a more detailed discussion of the performance of the SQD algorithm.
The SQD algorithm contains two phases. In Phase 1, the algorithm constructs a dense graph with at least $\frac{k}{2}$ nodes. In Phase 2, the algorithm tries to satisfy the cardinality constraint. If the dense graph found from the phase 1 contains more than $k$ nodes, the algorithm in phase $2$ removes nodes according to the $\sigma$-elimination order; otherwise, the algorithm randomly adds nodes.
We observe that SQD performs better than Greedy on the SNAP-DBLP dataset (see Table 2). This is because that on the SNAP-DBLP dataset, Phase 1 of SQD constructs a dense while small subgraph, and in the Phase 2 it only randomly adds extra nodes while still gaining larger density than Greedy.
We also observe that on some datasets SQD does not produce results. This is because the graph constructed after Phase 1 either does not have a $\sigma$-quasi-elimination ordering, or due to the high time and memory complexity of computing a $\sigma$-quasi-elimination ordering. For example, on small graphs, SBM-Balanced and SBM-DenseSubg in Table 2, and SBM-DenseSubg in Table 3, SQD constructs a graph with significantly larger than $10\%\cdot n_{0}$ nodes in Phase 1. In fact, the size of the graoh is equal to the size of the whole remaining graph.
Even for those datasets where SQD produces results, if computing the $\sigma$-quasi-elimination order is necessary, it is very costly. In Figure 12 on SBM-SparseSubg, we see that running SQD is even more costly than running SDP. We should also note that for the largest datasets (which are relatively sparse), such as SNAP-DBLP, SQD does compute the $\sigma$-elimination ordering and obtains results.
Appendix F Experiments for Max-Cut with $k$ refinements
In this section, we conduct an extensive evaluation of our algorithms on a variety of datasets to solve Max-Cut-kkR. Our focus in this section will be two primary questions:
RQ1:
Do our algorithms consistently produce a larger cut when initialized with a random partition?
RQ2:
Which algorithm is the best among our proposed algorithms?
F.1 Algorithms
We use our SDP and Greedy algorithms and denote them as SDP and Greedy, respectively.
Additionally, we utilize the following black-box solvers for Max-Cut to obtain algorithms via the reduction in Theorem 4.2:
  1. 1.
    A Greedy algorithm, where each node’s side is chosen in a way that maximizes the cut at each step. We denote it by B: Greedy.
  2. 2.
    A Local Search algorithm which makes refinements to each node’s side if the cut can be improved. The algorithm either starts with a randomly initialized partition or uses the result of the Greedy algorithm as the initial partition. We denote them as B: Local :I and B: Local :II, respectively.
  3. 3.
    The SDP algorithm of Goemans and Williamson [1995] which we denote by B: SDP.
F.2 Datasets
We use the same datasets as for DS-kkR. Since SBM-DenseSubg and SBM-SparseSubg essentially have the same graph structure and since in Max-Cut-kkR the ground-truth communities are not relevant for us, we use SBM-DenseSubg to represent them. The networks statistics are presented in Table 4.
Table 4: Network statistics, and average relative increase of cut in 55 runs with $k=50$. Here, $n$ and $m$ are the number of nodes and edges of the graph; $\mathsf{cut}_{0}$ is the average of 55 random cuts. Next, _ denotes that an algorithm does not finish in time (2 hours for SNAP datasets, and 30 minutes for other datasets).
dataset Network Property Algorithms Blackbox Methods
$n$ $m$ $\mathsf{cut}_{0}$ Greedy SDP B:Greedy B:SDP B:Local:I B:Local:II
SBM-Balanced 1 000 74 940 37 695.6 0.030 0.031 0.021 0.020 0.021 0.021
SBM-DenseSubg 1 000 81 531 40 761.8 0.027 0.028 0.019 0.018 0.020 0.019
Wiki-ES 205 372 185.2 0.438 0.533 0.421 0.462 0.386 0.426
Wiki-DE 768 3 059 1 531.6 0.157 0.182 0.118 0.142 0.126 0.126
Wiki-GB 2 168 18 617 9 314.6 0.070 0.074 0.035 0.049 0.043 0.038
Wiki-US 3 912 18 359 9 157.0 0.058 0.061 0.035 0.031 0.034 0.034
SNAP-DBLP 317 080 1 049 866 525 028.4 0.002 _ 0.001 _ 0.001 0.001
SNAP-Amazon 334 863 925 872 462 570.4 0.002 _ 0.001 _ 0.001 0.001
SNAP-Youtube 1 134 890 2 987 624 1 493 444.6 0.002 _ 0.000 _ _ -0.000
F.3 Evaluation
In our experiments, we proceeds as follows. We uniformly randomly partition each node in the graph to either side, and repeat this procedure $5$ times. At each time, we run our algorithms and calculate the cut. We compute the relative increase on the initial given cut at each run, and take the average of these $5$ runs. We present the results in Table 4.
Here, it is worth noting that for Max-Cut, a random partition already yields a 2-approximate solution compared to the maximum cut in expectation. Thus, since we initialize the original cut randomly, this means that the given cut will already be large (with a relatively large probability) and therefore improving the cut significantly will be difficult for our algorithms.
F.4 Performance
In Table 4, we report the average of the relative increases of the cut value by different algorithms. In Figure 15, we show the relative increase of the cut values for our Wikipedia politician networks for varying values of $k$.
We notice from the plots and the table that, for all the settings of $k$, SDP is always the best, as long as SDP outputs a result. As $k$ increases, there is a trend that SDP performs better. Even though Greedy does not perform better than SDP, in many cases, it performs better than the other competitors. This addresses RQ2.
Additionally, the local-search based algorithm B: Local :II only very slightly improves upon B: Greedy (which it uses to obtain its initial partition) and does not perform significantly better than B: Local :I (which uses a random initialization). This suggests that the initial solution returned by B: Greedy is not very good, as is also suggested by comparison with Greedy and SDP. However, we will see below that B: Greedy has a significantly lower running time than B: Local :I. This addresses RQ1.
Wiki-GB Wiki-DE SBM-Balanced SBM-DenseSubg
Wiki-US SNAP-DBLP SNAP-Amazon SNAP-Youtube
Figure 15: Relative increase of the cut value. We used $k\in[10,50]$ on smaller graphs, and $k\in[10,100]$ for larger graphs.
Running time dependency on $k$. In Figure 16 we present the running time on three datasets with varying values of $k$. We notice that SDP requires much longer time than B: SDP. This is because B: SDP only needs to solve the standard Max-Cut SDP, while SDP adds additional constraints, resulting in longer running time. B: Local :I requires more time than B: Local :II, this is because the local search algorithm with random initialization takes much longer time to converge.
Wiki-GB Wiki-DE SBM-Balanced SBM-DenseSubg
Wiki-US SNAP-DBLP SNAP-Amazon SNAP-Youtube
Figure 16: Running time of our algorithms. We set $k\in[10,50]$ on smaller graphs, and $k\in[10,100]$ for larger graphs.
Running time dependency on $n$. We increase the size of synthetic datasets SBM-DenseSubg and SBM-Balanced, and in Figure 17 we give the running time in terms of increasing $n$. SDP and B: SDP increase largely as $n$ increases, as we expect, indicating these two algorithms are not scalable.
(a) SBM-DenseSubg (b) SBM-Balanced
Figure 17: Running time of the algorithms for fixed $k=50$ and varying n=1 000,1 250,1 500,1 750,2 000n=1\,000,1\,250,1\,500,1\,750,2\,000.
F.5 A case study
Matakos et al. [2020] propose a method to maximize the diversity of a social network by flipping $k$ nodes’ opinions, where each opinion is either $1$ or 1-1. In our problem setting, the term diversity is mathematically equivalent to cut and the nodes of the graph can be partitioned based on their opinions (where the cut is given by all nodes with opinion $1$ on one side and all other nodes on the other side). It is important to note that our problem setting differs from theirs in several ways: they operate under a budget constraint, whereas we operate under a cardinality constraint; they require an inequality constraint, while our problem necessitates an equality constraint. We executed their algorithm on our datasets by setting the cost of all nodes to $1$, allowing for a comparison of our methods. The performance, running time, and the average number of changed nodes are presented in Table 5. We note that we do not report results for Wiki-US, since the algorithm of Matakos et al. [2020] runs out of memory on this dataset.
We observe that, somewhat surprisingly, the SDP algorithm proposed by Matakos et al. [2020] takes significantly more time than the SDP algorithm proposed in our paper (as shown in Figure 16). In addition, despite having the at most constraint instead of the equality constraint, their algorithm does not yield better performance. Specifically, on the dataset Wiki-DE, our SDP algorithm performs twice as well.
Table 5: The results of the SDP algorithm proposed by Matakos et al. [2020], where we set $k=50$ on several datasets. We use Matakos et al. to indicate the average relative increase of the cut value with their approach in 55 runs, and SDP to indicate the result of our SDP-based algorithm.
dataset SDP Matakos et al. Time (s) # changed nodes
Wiki-ES 0.533 0.408 22.1848 48.3
Wiki-DE 0.182 0.105 784.401 46.4
Wiki-GB 0.074 0.031 5780.306 48.5